### v1

# Holistic Optimization Framework for FPGA Accelerators

Abstract
Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization. Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential. High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives. However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management. While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space. Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
I Introduction
The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization. Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential. High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions. However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71]. Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance. Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions. Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design. This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs). In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks. This method increases parallelism and is particularly effective for computation-bound kernels. Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58]. However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained. While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks. This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out). Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks. Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use. A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus. This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation. The main contributions are:
  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture. This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
  • Automating our framework that is end-to-end comprehensive. It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14]. Importantly, it achieves these results without requiring DSE for HLS design generation.
II Background and Motivation
II-A HLS optimization
In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role. The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II). array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling. Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18]. On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions. For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers. The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
II-B Limitation of existing methods
To demonstrate the limitations of the existing framework and methodology we use 3mm. As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication. AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently. Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved. Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored. Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
1for (i = 0; i < 180; i++) // MM 1
2 for (j = 0; j < 190; j++) {
3 E[i][j] = 0.0; // S0
4 for (k = 0; k < 200; ++k)
5 E[i][j] += A[i][k] * B[k][j];} // S1
6for (i = 0; i < 190; i++) // MM 2
7 for (j = 0; j < 210; j++) {
8 F[i][j] = 0.0; // S2
9 for (k = 0; k < 220; ++k)
10 F[i][j] += C[i][k] * D[k][j];} // S3
11for (i = 0; i < 180; i++) // MM 3
12 for (j = 0; j < 210; j++) {
13 G[i][j] = 0.0; // S4
14 for (k = 0; k < 190; ++k)
15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 1: Code of 3mm
Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs. For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers. However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
As 3mm is computation-bound, increasing parallelism is vital. However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation. Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR. Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency. These gaps can lead to suboptimal performance, particularly in data-intensive scenarios. Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors. However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops. Expanding unroll factor options could better balance parallelism and performance.
For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74). These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method. However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
II-C Overview of Prometheus
We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently. Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel. The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs. Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1. We can see an example on Listing 3. Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded. Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
1/*************** Fused Task 0 ****************/
2float B[204][192]; load_B(B);
3for (i0 = 0; i0 < 18; i0++)//inter-tile loop
4 float A[10][204]; load_A(A);
5 for (j0 = 0; j0 < 6; j0++){//inter-tile loop
6 float E[10][32];
7 task_0(E); // S0 + intra-tile loops
8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
9#pragma HLS pipeline II=3
10 task_1(E, A, B); // S1 + intra-tile loops
11 store_E(E); sent_E(E);
12/*************** Fused Task 1 ****************/
13float C[190][222]; load_C(C);
14for (j0 = 0; j0 < 7; j0++)//inter-tile loop
15 float D[222][32]; load_D(D);
16 for (i0 = 0; i0 < 10; i0++){//inter-tile loop
17 float F[19][32];
18 task_2(F); // S2 + intra-tile loops
19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
20#pragma HLS pipeline II=3
21 task_3(F, C, D); // S3 + intra-tile loops
22 store_F(F); sent_F(F);
23/*************** Fused Task 2 ****************/
24float E[180][192];
25for (j0 = 0; j0 < 7; j0++)//inter-tile loop
26 float F[192][32];
27 receive_F(F);
28 for (i0 = 0; i0 < 18; i0++){//inter-tile loop
29 float G[10][32];
30 receive_E(E);
31 task_4(G); // S4 + intra-tile loops
32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
33#pragma HLS pipeline II=3
34 task_5(G, E, F); // S5 + intra-tile loops
35 store_G(G);
Listing 2: Pseudo Code of 3mm generated by Prometheus
Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks. Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready. These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels. Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
III Problem Formulation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline [15] each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
Prometheus’s solution space includes:
Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop. To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop. Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is infeasible, the inter-tile loop remains in its legal position. In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$. Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once. Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data. Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP, as described in Section IV. As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written. Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks  Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
IV NLP Formulation
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed by Pouget et al. [59, 58, 60], which we have adapted to address our specific requirements and constraints. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section III. We employ PoCC [54] to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC [69] then generates all legal permutations for each loop body.
Constant Description
$II_{l}$ II of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$TC^{l}_{ori}$ Original Trip Count of the loop $l$
$f_{a,l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$N_{FT}$ Number of fused task
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
$DSP$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning
$SLR$ Number of SLR available for the FPGA used
Variable Description
$TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
$TC^{l}$ Trip Count of the loop $l$ after padding
$S^{last}_{a}$ Size of the last dimension of the array $a$ transferred on-chip
$BW_{a}$ Bit width of the array $a$
$t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
$slr_{t}$ ID of the SLR use by the task $t$
Set Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
$\mathcal{L}_{s}$ The set of loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ The set of reduction loops which iterate the statement $s$
$\mathcal{L}_{a}^{last}$ The set of loop which iterate the last dimension of the array $a$
$\mathcal{L}_{inter},$ $\mathcal{L}_{intra}$ The set of loops which belong to the inter-tile and intra-tile respectively
$\mathcal{B}$ Set of possible burst size for the data type
$\mathcal{C}_{a_{d}}$ The set of loops which iterates the array $a$ at the dimension $d$
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
$\mathcal{P}_{s}$ All permutation of the loops which iterate the statement $s$
$\mathcal{F}_{i}$ The set of statements which belong to the fused task $i$
$\mathcal{T}$ The set of tasks in the dataflow
TABLE I: Overview of the constant, variable and set employed in formulating the NLP
IV-A Constraints
We now describe the constraints by using the code of Listings 1, 2 and 3.
1for (int j1 = 0; j1 < 32; j1++) {
2#pragma HLS unroll
3 for (int i1 = 0; i1 < 19; i1++) {
4#pragma HLS unroll
5 for (int k1 = 0; k1 < 3; k1++) {
6#pragma HLS unroll
7 j=j0*32+j1; i=i0*19+i1; k=k0*3+k1;
8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 3: Pseudo Code of the task_3
Data-tiling and Unroll Factor  The intra-tile transformation, as explained in Section III, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 3, for the array $F$, the loop $j$ (line 7 in Listing 1) has been split into $j0$ (line 14 in Listing 2) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
$\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$ (1)
$(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
Bit Width  $\mathcal{B}$ represents the possible bit width for the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
$\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement. Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
$\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
$\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$ (5)
Transfer and Reuse  Equation 6 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 7 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 2, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 7 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 2, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
$\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$ (6)
$\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$ (7)
On-chip Memory  Equation 8 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
$\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$ (8)
with $N_{a}$ being the number of double buffer for the array $a$.
Array Partitioning  Equation 9 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 10 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 3 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
$\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$ (9)
$\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$ (10)
DSP Utilization  Equation 11 constrains the number of DSPs used based on the available DSP resources. In contrast to [59, 58, 60], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously.
Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
$\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$ (11)
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12). Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
$\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (12)
IV-B Objective Function
The objective function models the dependency graph of the fused task. Independent tasks execute simultaneously, with latency set by the longest task. Dependent tasks overlap partially, with the second task starting once it receives enough data from the first. The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts. For our 3mm example, the graph is shown in Figure 1, and the corresponding latency is:
$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
T0T1T2T3T4T5
Figure 1: Dataflow graph of 3mm maximally distributed
The latency of the intra-task for a statement $s$ is determine similarly to [59, 58, 60] with Equation 13. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 14.
$Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$ (13)
$Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$ (14)
For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication. The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
$Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$ (15)
And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
V Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section IV gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
Communication with Off-Chip Memory  To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
Communication between the Fused Task  The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory. Data for these tasks resides on-chip. If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function. The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined. Padding is managed at the intra-tile level. For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements. In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
VI Evaluation
VI-A Setup and Experimental Evaluation
We evaluated our method using kernels from Polybench/C 4.2.1 [56] with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow [4]. This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We execute the designs for these frameworks using identical parameters as outlined in Section VI-A. For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [60], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [60], AutoDSE [67], ScaleHLS [79] (S-HLS), and Allo [14]. The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels. Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days.
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus [60], designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
Kernel Ours Sisyphus S-HLS Allo AutoDSE
2mm 308.38 195.09 37.13 46.58 0.41*
3mm 368.36 178.97 43.04 60.40* 1.74*
Atax 3.56 2.32 1.58 1.96 1.97*
Bicg 15.41 2.32 1.70 14.17 0.99*
Gemm 419.14 227.09 40.53 37.50 110.81*
Gesummv 10.21 2.28 1.78 8.85 1.98*
Mvt 14.65 5.54 7.39 8.77 7.80*
Symm 212.20 200.30 0.06 16.72 14.68*
Syr2k 133.77 75.01 0.04 20.59 12.35*
Syrk 316.25 211.00 0.54 41.11 23.16*
Trmm 193.47 166.72 0.07 5.10 0.02*
PI (Avg) 2.39x 927.20x 8.58x 973.14
PI (gmean) 2.03x 48.03x 4.92x 25.82
TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar. Allo retains the original code structure by permuting the reduction loop outermost. The non-reduction loop is fully unrolled, and the reduction loop is pipelined. Prometheus partially unrolls both loops and pipelines the reduction loop.
On Board Evaluation  Table III presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
For Bicg and Atax, targeting 60% of resources on one SLR caused congestion, requiring regeneration with a 55% constraint. The 3mm bitstream from AutoDSE succeeded only with a 15% constraint.
Kernel T (ms) GF/s DSP BRAM L(K) FF(K) F (MHz)
1 SLR Sisyphus 2mm 1.20 30.57 556 510 213 276 220
3mm 1.52 29.89 984 611 230 300 220
Atax 0.62 1.03 173 450 240 250 220
Bicg 0.63 1.02 173 451 238 265 217
1 SLR AutoDSE 2mm 92.25 0.40 963 353.5 287 292 205
3mm 110.34 0.41 1117 470 278 306 220
Atax 2.88 0.22 452 630.5 170 212 220
Bicg 1.13 0.56 196 867.5 168 217 214
1 SLR Ours 2mm 0.56 65.13 1941 635.5 371 454 216
3mm 0.87 51.95 1551 635.5 342 423 220
Atax 0.24 2.62 1081 533.5 234 287 184
Bicg 0.15 4.04 732 311.5 250 302 220
3 SLR Ours 2mm 0.29 125.54 2752 546 428 549 220
3mm 0.34 134.07 4379 600 684 840 207
Atax 0.20 3.10 1823 634.5 405 539 137
Bicg 0.14 4.34 1226 241 291 380 177
TABLE III: On-board evaluation comparison
Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [56] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
Permutations are evident in the implementations of 3mm, Atax, and Bicg. For 2mm and 3mm, the NLP opted to fully transfer array B instead of overlapping computation and load, as the load operation had higher latency than computation. In 2mm, the NLP retains the original loop order. Array tmp is transferred from the first task to the second, with both tasks iterating over the first dimension using loop $i$. This enables a $4\times 32$ data tile to be sent to the second task, which starts computation once a $4\times 192$ tile of tmp(FT1) is filled.
Fused Statement Loop   Order Data Tile Size
2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
3mm FT0: S0, S1, FT1: S2, S3, FT2: S4, S5 FT0: i,j,k, FT1: j,i,k FT2: j,i,k E(FT0): $9\times 16$, A: $9\times 200$, B: $200\times 192$, F(FT1): $10\times 10$, C: $10\times 224$, D: $224\times 10$, G: $9\times 10$, E(FT2): $9\times 192$, F(FT2): $192\times 10$,
Atax FT0: S1, S2, FT1: S0, S3 FT0: i,j, FT1: j,i tmp(FT0): $56$, y: $16$, A(FT0): $392\times 416$, A(FT1): $392\times 16$, tmp(FT1): $392$
Bicg FT0: S1, S2, FT1: S0, S3 FT0: j,i FT1: i,j s: $16$, A (FT0): $416\times 16$, r: $416$, q: $10$, A (FT1): $10\times 400$, p: $400$,
TABLE IV: Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table III for 1 SLR
As other frameworks cannot generate bitstreams for 3 SLRs without human intervention, we compare results for 1 and 3 SLRs using our framework. For 2mm and 3mm, performance improves due to increased resource utilization. However, for atax and bicg, where the bottleneck is memory transfer between off-chip and on-chip rather than parallelism, the improvement is negligible.
VII Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [22], synchronous dataflow languages [42, 11], and for FPGA applications [53, 12, 3, 25, 10, 5, 75, 14]. DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [14, 75], built on HeteroCL [38], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [71, 20, 33, 39, 9] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [58, 59, 60], bottleneck DSE [67], and GNN-based latency and resource estimation [66, 64, 65, 8, 68, 73, 74]. However, these approaches lack effective integration of dataflow optimization techniques.
Code Transformation  Code transformation has been explored for CPUs [13, 55, 37, 7], GPUs [70], and FPGAs [57, 84, 83, 44, 48, 47, 46, 17]. Pluto [13] minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations [57] leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works [84, 83] selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While [44, 48, 47, 46, 17] focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The [79, 38, 31, 14, 78] compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper [60] described in Section II generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works [49, 45] use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling [45] focus on CPUs, while Wedler [61] optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses [30, 52] and improving memory transfers [72], but its use for varying unroll factors on FPGAs remains underexplored.
VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.

### v2

@@
 Abstract
-Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization.
-Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential.
-High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives.
-However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management.
-While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
-This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space.
-Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
+Customized accelerators have transformed modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization.
+Field-Programmable Gate Arrays play a critical role in this paradigm, offering unmatched flexibility and high-performance potential.
+High-Level Synthesis and source-to-source compilers have simplified Field-Programmable Gate Arrays design by translating high-level programming languages into hardware descriptions enriched with directives.
+However, achieving high Quality of Results remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
+This paper presents Prometheus, a comprehensive framework that integrates key optimizations, including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution, into a unified design space.
+By leveraging Non-Linear Problem methodologies, Prometheus explores this design space to deliver efficient solutions under strict resource constraints, facilitating automatic bitstream generation.
+The framework addresses limitations in existing methods and demonstrates superior performance, achieving high Quality of Results across complex computation-bound applications.
 I Introduction
-The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization.
-Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential.
-High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions.
-However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
-Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71].
-Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance.
-Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions.
-Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design.
-This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs).
-In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
-Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks.
-This method increases parallelism and is particularly effective for computation-bound kernels.
-Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58].
-However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained.
-While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
-Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks.
-This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out).
-Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks.
-Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
-Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use.
-A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
-Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus.
-This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation.
-The main contributions are:
-  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture.
-This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
-  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
-  • Automating our framework that is end-to-end comprehensive.
-It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
-  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14].
-Importantly, it achieves these results without requiring DSE for HLS design generation.
+The rise of customized accelerators has transformed modern computing, delivering substantial improvements in energy efficiency and performance through hardware specialization.
+Field-Programmable Gate Arrays (FPGAs) have emerged as a particularly versatile solution, valued for their flexibility and ability to achieve high performance across diverse applications.
+High-level synthesis (HLS) tools and source-to-source compilers [20, 80, 75, 32, 61, 49, 76, 31, 78, 37, 79, 42, 59, 9] have simplified hardware design by translating high-level programming languages (e.g., C++, Python) enriched with hardware-specific directives into synthesizable hardware descriptions.
+However, the effectiveness of these tools depends heavily on the quality of the input code and the strategic application of hardware directives, making it challenging to achieve high Quality of Results (QoR) without significant manual effort.
+Despite advancements in the field, designing optimized FPGA solutions remains a complex task.
+Existing approaches [57, 58, 66, 65, 63, 25, 81, 65, 37, 78, 77, 59, 70, 9, 31] often address specific aspects of FPGA design in isolation, leaving gaps in areas such as efficient code transformation, effective concurrency management, and resource optimization.
+Moreover, challenges such as overlapping computation and communication, balancing intra-task parallelism, and managing FPGA resource constraints—including multi-Super Logic Region (SLR) architectures—further complicate the development process.
+These challenges are exacerbated by the lack of holistic frameworks that integrate multiple optimization strategies while ensuring synthesizable output.
+Two commonly used paradigms—shared buffer and dataflow—have shown promise in improving FPGA performance.
+Shared buffers enable data reuse and promote parallelism by caching information on-chip, which is particularly beneficial for computation-bound kernels.
+However, their efficacy is often constrained by the limited on-chip memory, preventing full data caching and hindering the ability to overlap computation and communication effectively.
+On the other hand, the dataflow paradigm facilitates efficient data movement and task concurrency by leveraging FIFO (First-In-First-Out) communication.
+While this approach is effective for memory-bound kernels, it struggles to exploit shared memory and parallelism within tasks, and its independent task compilation can limit resource sharing opportunities.
+This work addresses these challenges by focusing on optimizing HLS for FPGA design, with a particular emphasis on reducing latency while adhering to resource constraints.
+The goal is to transform affine high-level code (e.g., Listing 1) into an SLR-aware HLS-C++ design that achieves high QoR and facilitates automated bitstream generation.
+Achieving this requires addressing multiple challenges simultaneously, including task-to-SLR mapping, balancing computation and communication, and efficiently regenerating code to adapt to resource conflicts or spatial congestion.
+Existing methods fall short of providing an integrated solution capable of achieving these objectives, highlighting the need for a unified approach.
+To address these limitations, we propose Prometheus, an integrated framework that combines the benefits of shared buffer and dataflow paradigms to optimize HLS-based FPGA design.
+The key contributions of this work include:
+  • A unified design space supporting tiled dataflow architectures, task fusion, tiling strategies, loop permutation, overlapping computation and communication, and concurrent task execution.
+  • The application of a Non-Linear Problem (NLP) formulation to discover optimal design parameters within strict resource constraints, enabling efficient bitstream generation.
+  • An end-to-end automated framework that generates OpenCL host code for correctness verification, performs Vitis HLS simulations, and produces the final bitstream.
+  • Rigorous evaluation through RTL simulation and on-board evaluation, demonstrating superior QoR compared to state-of-the-art tools like AutoDSE [66], Sisyphus [59], and ScaleHLS [78].
 II Background and Motivation
 II-A HLS Optimization
-In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role.
-The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II).
-array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling.
-Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18].
-On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions.
-For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers.
-The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
-Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation.
-Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
-II-B Limitation of existing methods
-To demonstrate the limitations of the existing framework and methodology we use 3mm.
-As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
-Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication.
-AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
-Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently.
-Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved.
-Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
-Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored.
-Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
+HLS optimizations rely heavily on directives such as unroll, pipeline, and array_partitioning, which play a pivotal role in enhancing performance.
+The unroll pragma duplicates the loop body in the RTL design to enable parallel execution of loop iterations, while pipeline overlaps loop operations, reducing latency based on the initiation interval (II).
+Similarly, array_partitioning divides arrays into on-chip memory banks, allowing simultaneous access and facilitating more effective unrolling.
+Advanced optimizations also leverage shared buffers to reduce off-chip memory transfers, ping-pong buffering to overlap data transfer and computation, and coarse-grained parallelism through duplicated processing elements [19].
+Modern AMD/Xilinx FPGAs support data transfers with bit widths up to 512 bits, determined by problem dimensions.
+For instance, transferring 216 floats using a 256-bit width (8 floats per transfer) requires 27 transfers.
+HLS tools like Vitis HLS support the dataflow pragma, which organizes computations into independent actors communicating through FIFO queues [40].
+This structure enables concurrent execution by allowing each actor to process inputs and generate outputs as soon as data becomes available.
+Performance evaluations often combine estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation.
+While estimations and simulations avoid place-and-route overheads, on-board evaluations impose stricter constraints for generating deployable bitstreams.
+II-B Limitations of Existing Methods
+The limitations of current HLS optimization frameworks are illustrated through the 3mm kernel, which performs two independent matrix multiplications followed by a dependent multiplication.
+As shown in Listing 1, each matrix multiplication involves nested loops that can benefit from parallelism, unrolling, and communication-computation overlap.
 1for (i = 0; i < 180; i++) // MM 1
@@
 Listing 1: Code of 3mm
-Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs.
-For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers.
-However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
-As 3mm is computation-bound, increasing parallelism is vital.
-However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation.
-Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
-Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR.
-Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
-Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency.
-These gaps can lead to suboptimal performance, particularly in data-intensive scenarios.
-Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors.
-However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops.
-Expanding unroll factor options could better balance parallelism and performance.
-For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74).
-These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method.
-However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
+Shared Buffer  Frameworks such as AutoDSE [66], Sisyphus [59], ScaleHLS [78], and others [65, 63, 58, 57] optimize 3mm primarily through sequential execution of matrix multiplications.
+This approach enables reuse of DSP resources but fails to leverage opportunities for concurrent execution, such as running the first two multiplications in parallel.
+Current frameworks transfer arrays at different loop levels to manage on-chip memory but lack automatic mechanisms to overlap computation with communication effectively.
+Although shared buffering could support concurrent execution of independent tasks, its application remains limited.
+Extending concurrency to all three multiplications would require task fusion, which may constrain compiler optimizations.
+Dataflow  The dataflow paradigm, as implemented in frameworks like StreamHLS [9], divides computations into independent tasks, with communication managed through FIFOs.
+For the 3mm kernel in Listing 1, dataflow enables concurrent execution of the first two matrix multiplications, while the third begins as soon as its inputs are ready.
+This approach overlaps computation and communication, effectively managing task dependencies.
+As 3mm is computation-bound, increasing parallelism is essential.
+However, parallelism in pure dataflow methods is constrained by the FIFO bit width of the receiving task and the number of FIFOs.
+Increasing FIFO bit widths or adding more FIFOs to handle higher concurrency often results in routing congestion, making it challenging to generate a valid bitstream.
+Additionally, the lack of shared memory prevents data reuse across tasks, leading to repeated off-chip transfers or redundant buffers.
+Mixed Methods  Mixed approaches, such as those used by Allo [15] and HIDA [77], combine shared buffering and dataflow paradigms.
+Allo extends HeteroCL [37] within the MLIR framework [39], supporting code transformation and dataflow scheduling.
+However, these methods lack automatic data-tile selection and pragma insertion, limiting their ability to optimize QoR effectively.
+General Limitations  Most existing HLS frameworks fail to address key challenges such as automatic ping-pong buffering, tile size adjustments, and padding to maximize bit widths for communication efficiency.
+Unroll factors are often restricted to evenly divide problem dimensions, which can lead to inefficiencies with partial tiles.
+Expanding these factors and incorporating dynamic adjustments could significantly enhance parallelism and performance.
 II-C Overview of Prometheus
-We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently.
-Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
-The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel.
-The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs.
-Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1.
-We can see an example on Listing 3.
-Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded.
-Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
-It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
+We present Prometheus, a unified optimization framework designed to efficiently explore design spaces using a Non-Linear Problem (NLP) cost model.
+Prometheus simplifies the development of FPGA designs by offering an engineer-friendly interface with adjustable parameters such as tile size, array partitioning, and unroll factors.
+It is SLR-aware, ensuring effective task distribution across Super Logic Regions (SLRs), which simplifies bitstream generation while optimizing the balance between performance and resource utilization.
+The pseudo-code in Listing 2 demonstrates how Prometheus processes the 3mm kernel.
+Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
+Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 1. An example of such a task is illustrated in Listing 3.
+Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
+Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 1), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
+The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
 1/*************** Fused Task 0 ****************/
@@
 6 float E[10][32];
-7 task_0(E); // S0 + intra-tile loops
+7 task0(E); // S0 + intra-tile loops
 8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
 9#pragma HLS pipeline II=3
-10 task_1(E, A, B); // S1 + intra-tile loops
+10 task1(E, A, B); // S1 + intra-tile loops
 11 store_E(E); sent_E(E);
@@
 17 float F[19][32];
-18 task_2(F); // S2 + intra-tile loops
+18 task2(F); // S2 + intra-tile loops
 19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
 20#pragma HLS pipeline II=3
-21 task_3(F, C, D); // S3 + intra-tile loops
+21 task3(F, C, D); // S3 + intra-tile loops
 22 store_F(F); sent_F(F);
@@
 30 receive_E(E);
-31 task_4(G); // S4 + intra-tile loops
+31 task4(G); // S4 + intra-tile loops
 32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
 33#pragma HLS pipeline II=3
-34 task_5(G, E, F); // S5 + intra-tile loops
+34 task5(G, E, F); // S5 + intra-tile loops
 35 store_G(G);
 Listing 2: Pseudo Code of 3mm generated by Prometheus
-Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks.
-Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready.
-These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels.
-Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
-III Problem Formulation
+Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance.
+Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
+For the 3mm kernel optimized with RTL simulation, Prometheus achieved a throughput of 368.36 GF/s, significantly surpassing state-of-the-art frameworks such as Sisyphus (178.97 GF/s), Allo (60.40 GF/s), ScaleHLS (43.04 GF/s), and AutoDSE (1.74 GF/s).
+These results demonstrate the effectiveness of Prometheus in overcoming the limitations of existing methods and achieving superior QoR.
+III Code Transformation
 In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time.
@@
 To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
-Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
+Dependency Graph Creation  Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
 ISCC [68] verifies the legality of these transformations, ensuring dependencies are preserved.
-After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies.
+After achieving full distribution, we construct a dependency graph using PoCC [53].
+PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph.
+In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies.
 Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties.
+This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once.
 If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
 Prometheus’s solution space includes:
-Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes.
+Data-tiling and Padding  In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes.
 Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets.
 Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied.
-If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop.
+For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
 To optimize memory transfers, padding is applied to arrays in two ways.
@@
 Tile sizes are consistent within a fused task but vary between tasks.
-For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
-Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop.
-Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner).
-If permutation is infeasible, the inter-tile loop remains in its legal position.
-In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism.
+In Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
+Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops.
+Using ISCC [68], we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner).
+If permutation is not feasible, the inter-tile loop retains its legal position.
+For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction.
+For instance, in Listing 2, the inter-tile loops on lines 3 and 5 iterate over task0 and task1.
+The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip.
 Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access.
@@
 For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$.
-Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once.
-Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data.
-Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
 Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile.
@@
 This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops.
-This order will subsequently be determined by the NLP, as described in Section IV.
+This order will subsequently be determined by the NLP (cf.
+Section IV).
 As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
@@
 If defined or transferred before loops, $l=0$.
-Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written.
+Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written.
 Following [19], we perform an initial load, then overlap loading the next tile with computing the current one.
@@
 Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic.
-For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
+For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
 IV NLP Formulation
@@
 ISCC [68] then generates all legal permutations for each loop body.
-Constant Description
+Table I delineates the sets, variables, and constants utilized in our NLP formulation.
+Constants Description
 $II_{l}$ II of the loop $l$
@@
 $SLR$ Number of SLR available for the FPGA used
-Variable Description
+Variables Description
 $TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
@@
 $t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
-$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
+$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
 $slr_{t}$ ID of the SLR use by the task $t$
-Set Description
+Sets Description
 $\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
@@
 $(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
-Bit Width  $\mathcal{B}$ represents the possible bit width for the data type.
+Bit Width  $\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type.
 Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$.
@@
 $\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
-Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement.
-Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task.
+Permutation  Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task.
 For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$).
 However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
-$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
 $\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
@@
 SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 11).
-Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
+Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
 $\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (11)
@@
 Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
-Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory.
-Data for these tasks resides on-chip.
-If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function.
-The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined.
-Padding is managed at the intra-tile level.
-For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements.
-In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
+Intra-Task  Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory.
+Data for these tasks resides entirely on-chip.
+If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined.
+While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
+Padding is handled at the intra-tile level.
+Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage.
+For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
 Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering.
 Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity.
-Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required.
+Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required.
 This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
-Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
-SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
+Concurrent Execution  Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites.
+Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
+SLR Management  The NLP determines the SLR ID for each task, specifying where it will be executed.
+Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
 VI Evaluation
@@
 The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels.
-To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE.
-We execute the designs for these frameworks using identical parameters as outlined in Section VI-A.
+To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE.
+We do not include a comparison with Stream-HLS [9] due to the unavailability of its source code.
 For ScaleHLS [78] and Allo [15], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory.
@@
 VI-B Comparison
-The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
+The average time to solve the NLP and generate the designs presented in this section is 31.70 seconds, with a maximum of 178.41 seconds for 3mm with 3 SLRs.
 RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [59], AutoDSE [66], ScaleHLS [78] (S-HLS), and Allo [15].
@@
 PI (gmean) 2.03x 48.03x 4.92x 25.82
-TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
+TABLE II: Comparison of the throughput (GF/s) with RTL Simulation
 Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar.
@@
 $Si$ represent the statement in position $i$ in the code.
-The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
+The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
 Permutations are evident in the implementations of 3mm, Atax, and Bicg.
@@
 DaCe [11] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention.
+Stream-HLS [9] automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory.
+Additionally, it offers very limited opportunities for parallelism.
 Flower [5] automates FPGA dataflow development but limits parallelism.
@@
 Application-specific frameworks [17, 29, 50, 22, 62, 24, 35, 6] demonstrate dataflow advantages but do not generalize across domains.
+RapidStream-TAPA [27, 28] enhances the performance of dataflow designs and automates SLR placement.
+However, it requires an optimized kernel with a dataflow structure as input.
 Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [57, 58, 59], bottleneck DSE [66], and GNN-based latency and resource estimation [65, 63, 64, 8, 67, 72, 73].
@@
 In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework.
-By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints.
+By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to minimize latency within stringent resource constraints.
 Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.

### v3

@@
 Abstract
-Customized accelerators have transformed modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization.
-Field-Programmable Gate Arrays play a critical role in this paradigm, offering unmatched flexibility and high-performance potential.
-High-Level Synthesis and source-to-source compilers have simplified Field-Programmable Gate Arrays design by translating high-level programming languages into hardware descriptions enriched with directives.
-However, achieving high Quality of Results remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
-This paper presents Prometheus, a comprehensive framework that integrates key optimizations, including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution, into a unified design space.
-By leveraging Non-Linear Problem methodologies, Prometheus explores this design space to deliver efficient solutions under strict resource constraints, facilitating automatic bitstream generation.
-The framework addresses limitations in existing methods and demonstrates superior performance, achieving high Quality of Results across complex computation-bound applications.
+Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization.
+Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential.
+High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives.
+However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management.
+While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
+This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space.
+Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
 I Introduction
-The rise of customized accelerators has transformed modern computing, delivering substantial improvements in energy efficiency and performance through hardware specialization.
-Field-Programmable Gate Arrays (FPGAs) have emerged as a particularly versatile solution, valued for their flexibility and ability to achieve high performance across diverse applications.
-High-level synthesis (HLS) tools and source-to-source compilers [20, 80, 75, 32, 61, 49, 76, 31, 78, 37, 79, 42, 59, 9] have simplified hardware design by translating high-level programming languages (e.g., C++, Python) enriched with hardware-specific directives into synthesizable hardware descriptions.
-However, the effectiveness of these tools depends heavily on the quality of the input code and the strategic application of hardware directives, making it challenging to achieve high Quality of Results (QoR) without significant manual effort.
-Despite advancements in the field, designing optimized FPGA solutions remains a complex task.
-Existing approaches [57, 58, 66, 65, 63, 25, 81, 65, 37, 78, 77, 59, 70, 9, 31] often address specific aspects of FPGA design in isolation, leaving gaps in areas such as efficient code transformation, effective concurrency management, and resource optimization.
-Moreover, challenges such as overlapping computation and communication, balancing intra-task parallelism, and managing FPGA resource constraints—including multi-Super Logic Region (SLR) architectures—further complicate the development process.
-These challenges are exacerbated by the lack of holistic frameworks that integrate multiple optimization strategies while ensuring synthesizable output.
-Two commonly used paradigms—shared buffer and dataflow—have shown promise in improving FPGA performance.
-Shared buffers enable data reuse and promote parallelism by caching information on-chip, which is particularly beneficial for computation-bound kernels.
-However, their efficacy is often constrained by the limited on-chip memory, preventing full data caching and hindering the ability to overlap computation and communication effectively.
-On the other hand, the dataflow paradigm facilitates efficient data movement and task concurrency by leveraging FIFO (First-In-First-Out) communication.
-While this approach is effective for memory-bound kernels, it struggles to exploit shared memory and parallelism within tasks, and its independent task compilation can limit resource sharing opportunities.
-This work addresses these challenges by focusing on optimizing HLS for FPGA design, with a particular emphasis on reducing latency while adhering to resource constraints.
-The goal is to transform affine high-level code (e.g., Listing 1) into an SLR-aware HLS-C++ design that achieves high QoR and facilitates automated bitstream generation.
-Achieving this requires addressing multiple challenges simultaneously, including task-to-SLR mapping, balancing computation and communication, and efficiently regenerating code to adapt to resource conflicts or spatial congestion.
-Existing methods fall short of providing an integrated solution capable of achieving these objectives, highlighting the need for a unified approach.
-To address these limitations, we propose Prometheus, an integrated framework that combines the benefits of shared buffer and dataflow paradigms to optimize HLS-based FPGA design.
-The key contributions of this work include:
-  • A unified design space supporting tiled dataflow architectures, task fusion, tiling strategies, loop permutation, overlapping computation and communication, and concurrent task execution.
-  • The application of a Non-Linear Problem (NLP) formulation to discover optimal design parameters within strict resource constraints, enabling efficient bitstream generation.
-  • An end-to-end automated framework that generates OpenCL host code for correctness verification, performs Vitis HLS simulations, and produces the final bitstream.
-  • Rigorous evaluation through RTL simulation and on-board evaluation, demonstrating superior QoR compared to state-of-the-art tools like AutoDSE [66], Sisyphus [59], and ScaleHLS [78].
+The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization.
+Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential.
+High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions.
+However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
+Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71].
+Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance.
+Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions.
+Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design.
+This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs).
+In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
+Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks.
+This method increases parallelism and is particularly effective for computation-bound kernels.
+Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58].
+However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained.
+While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
+Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks.
+This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out).
+Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks.
+Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
+Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use.
+A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
+Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus.
+This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation.
+The main contributions are:
+  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture.
+This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
+  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
+  • Automating our framework that is end-to-end comprehensive.
+It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
+  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14].
+Importantly, it achieves these results without requiring DSE for HLS design generation.
 II Background and Motivation
 II-A HLS optimization
-HLS optimizations rely heavily on directives such as unroll, pipeline, and array_partitioning, which play a pivotal role in enhancing performance.
-The unroll pragma duplicates the loop body in the RTL design to enable parallel execution of loop iterations, while pipeline overlaps loop operations, reducing latency based on the initiation interval (II).
-Similarly, array_partitioning divides arrays into on-chip memory banks, allowing simultaneous access and facilitating more effective unrolling.
-Advanced optimizations also leverage shared buffers to reduce off-chip memory transfers, ping-pong buffering to overlap data transfer and computation, and coarse-grained parallelism through duplicated processing elements [19].
-Modern AMD/Xilinx FPGAs support data transfers with bit widths up to 512 bits, determined by problem dimensions.
-For instance, transferring 216 floats using a 256-bit width (8 floats per transfer) requires 27 transfers.
-HLS tools like Vitis HLS support the dataflow pragma, which organizes computations into independent actors communicating through FIFO queues [40].
-This structure enables concurrent execution by allowing each actor to process inputs and generate outputs as soon as data becomes available.
-Performance evaluations often combine estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation.
-While estimations and simulations avoid place-and-route overheads, on-board evaluations impose stricter constraints for generating deployable bitstreams.
-II-B Limitations of Existing Methods
-The limitations of current HLS optimization frameworks are illustrated through the 3mm kernel, which performs two independent matrix multiplications followed by a dependent multiplication.
-As shown in Listing 1, each matrix multiplication involves nested loops that can benefit from parallelism, unrolling, and communication-computation overlap.
+In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role.
+The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II).
+array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling.
+Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18].
+On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions.
+For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers.
+The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
+Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation.
+Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
+II-B Limitation of existing methods
+To demonstrate the limitations of the existing framework and methodology we use 3mm.
+As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
+Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication.
+AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
+Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently.
+Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved.
+Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
+Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored.
+Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
 1for (i = 0; i < 180; i++) // MM 1
@@
 Listing 1: Code of 3mm
-Shared Buffer  Frameworks such as AutoDSE [66], Sisyphus [59], ScaleHLS [78], and others [65, 63, 58, 57] optimize 3mm primarily through sequential execution of matrix multiplications.
-This approach enables reuse of DSP resources but fails to leverage opportunities for concurrent execution, such as running the first two multiplications in parallel.
-Current frameworks transfer arrays at different loop levels to manage on-chip memory but lack automatic mechanisms to overlap computation with communication effectively.
-Although shared buffering could support concurrent execution of independent tasks, its application remains limited.
-Extending concurrency to all three multiplications would require task fusion, which may constrain compiler optimizations.
-Dataflow  The dataflow paradigm, as implemented in frameworks like StreamHLS [9], divides computations into independent tasks, with communication managed through FIFOs.
-For the 3mm kernel in Listing 1, dataflow enables concurrent execution of the first two matrix multiplications, while the third begins as soon as its inputs are ready.
-This approach overlaps computation and communication, effectively managing task dependencies.
-As 3mm is computation-bound, increasing parallelism is essential.
-However, parallelism in pure dataflow methods is constrained by the FIFO bit width of the receiving task and the number of FIFOs.
-Increasing FIFO bit widths or adding more FIFOs to handle higher concurrency often results in routing congestion, making it challenging to generate a valid bitstream.
-Additionally, the lack of shared memory prevents data reuse across tasks, leading to repeated off-chip transfers or redundant buffers.
-Mixed Methods  Mixed approaches, such as those used by Allo [15] and HIDA [77], combine shared buffering and dataflow paradigms.
-Allo extends HeteroCL [37] within the MLIR framework [39], supporting code transformation and dataflow scheduling.
-However, these methods lack automatic data-tile selection and pragma insertion, limiting their ability to optimize QoR effectively.
-General Limitations  Most existing HLS frameworks fail to address key challenges such as automatic ping-pong buffering, tile size adjustments, and padding to maximize bit widths for communication efficiency.
-Unroll factors are often restricted to evenly divide problem dimensions, which can lead to inefficiencies with partial tiles.
-Expanding these factors and incorporating dynamic adjustments could significantly enhance parallelism and performance.
+Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs.
+For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers.
+However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
+As 3mm is computation-bound, increasing parallelism is vital.
+However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation.
+Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
+Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR.
+Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
+Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency.
+These gaps can lead to suboptimal performance, particularly in data-intensive scenarios.
+Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors.
+However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops.
+Expanding unroll factor options could better balance parallelism and performance.
+For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74).
+These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method.
+However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
 II-C Overview of Prometheus
-We present Prometheus, a unified optimization framework designed to efficiently explore design spaces using a Non-Linear Problem (NLP) cost model.
-Prometheus simplifies the development of FPGA designs by offering an engineer-friendly interface with adjustable parameters such as tile size, array partitioning, and unroll factors.
-It is SLR-aware, ensuring effective task distribution across Super Logic Regions (SLRs), which simplifies bitstream generation while optimizing the balance between performance and resource utilization.
-The pseudo-code in Listing 2 demonstrates how Prometheus processes the 3mm kernel.
-Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
-Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 1. An example of such a task is illustrated in Listing 3.
-Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
-Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 1), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
-The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
+We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently.
+Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
+The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel.
+The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs.
+Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1.
+We can see an example on Listing 3.
+Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded.
+Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
+It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
 1/*************** Fused Task 0 ****************/
@@
 6 float E[10][32];
-7 task0(E); // S0 + intra-tile loops
+7 task_0(E); // S0 + intra-tile loops
 8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
 9#pragma HLS pipeline II=3
-10 task1(E, A, B); // S1 + intra-tile loops
+10 task_1(E, A, B); // S1 + intra-tile loops
 11 store_E(E); sent_E(E);
@@
 17 float F[19][32];
-18 task2(F); // S2 + intra-tile loops
+18 task_2(F); // S2 + intra-tile loops
 19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
 20#pragma HLS pipeline II=3
-21 task3(F, C, D); // S3 + intra-tile loops
+21 task_3(F, C, D); // S3 + intra-tile loops
 22 store_F(F); sent_F(F);
@@
 30 receive_E(E);
-31 task4(G); // S4 + intra-tile loops
+31 task_4(G); // S4 + intra-tile loops
 32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
 33#pragma HLS pipeline II=3
-34 task5(G, E, F); // S5 + intra-tile loops
+34 task_5(G, E, F); // S5 + intra-tile loops
 35 store_G(G);
 Listing 2: Pseudo Code of 3mm generated by Prometheus
-Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance.
-Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
-For the 3mm kernel optimized with RTL simulation, Prometheus achieved a throughput of 368.36 GF/s, significantly surpassing state-of-the-art frameworks such as Sisyphus (178.97 GF/s), Allo (60.40 GF/s), ScaleHLS (43.04 GF/s), and AutoDSE (1.74 GF/s).
-These results demonstrate the effectiveness of Prometheus in overcoming the limitations of existing methods and achieving superior QoR.
-III Code Transformation
+Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks.
+Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready.
+These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels.
+Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
+III Problem Formulation
 In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time.
@@
 To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
-Dependency Graph Creation  Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
+Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
 ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved.
-After achieving full distribution, we construct a dependency graph using PoCC [53].
-PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph.
-In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies.
+After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies.
 Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties.
-This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once.
 If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
 Prometheus’s solution space includes:
-Data-tiling and Padding  In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes.
+Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes.
 Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets.
 Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied.
-For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
+If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop.
 To optimize memory transfers, padding is applied to arrays in two ways.
@@
 Tile sizes are consistent within a fused task but vary between tasks.
-In Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
-Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops.
-Using ISCC [68], we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner).
-If permutation is not feasible, the inter-tile loop retains its legal position.
-For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction.
-For instance, in Listing 2, the inter-tile loops on lines 3 and 5 iterate over task0 and task1.
-The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip.
+For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
+Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop.
+Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner).
+If permutation is infeasible, the inter-tile loop remains in its legal position.
+In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism.
 Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access.
@@
 For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$.
+Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once.
+Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data.
+Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
 Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile.
@@
 This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops.
-This order will subsequently be determined by the NLP (cf.
-Section IV).
+This order will subsequently be determined by the NLP, as described in Section IV.
 As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
@@
 If defined or transferred before loops, $l=0$.
-Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written.
+Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written.
 Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
@@
 Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic.
-For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
+For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
 IV NLP Formulation
@@
 ISCC [69] then generates all legal permutations for each loop body.
-Table I delineates the sets, variables, and constants utilized in our NLP formulation.
-Constants Description
+Constant Description
 $II_{l}$ II of the loop $l$
@@
 $SLR$ Number of SLR available for the FPGA used
-Variables Description
+Variable Description
 $TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
@@
 $t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
-$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
+$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
 $slr_{t}$ ID of the SLR use by the task $t$
-Sets Description
+Set Description
 $\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
@@
 $(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
-Bit Width  $\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type.
+Bit Width  $\mathcal{B}$ represents the possible bit width for the data type.
 Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$.
@@
 $\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
-Permutation  Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task.
+Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement.
+Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task.
 For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$).
 However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
+$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
 $\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
@@
 SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12).
-Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
+Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
 $\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (12)
@@
 Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
-Intra-Task  Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory.
-Data for these tasks resides entirely on-chip.
-If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined.
-While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
-Padding is handled at the intra-tile level.
-Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage.
-For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
+Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory.
+Data for these tasks resides on-chip.
+If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function.
+The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined.
+Padding is managed at the intra-tile level.
+For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements.
+In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
 Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering.
 Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity.
-Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required.
+Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required.
 This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
-Concurrent Execution  Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites.
-Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
-SLR Management  The NLP determines the SLR ID for each task, specifying where it will be executed.
-Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
+Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
+SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
 VI Evaluation
@@
 The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels.
-To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE.
-We do not include a comparison with Stream-HLS [9] due to the unavailability of its source code.
+To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE.
+We execute the designs for these frameworks using identical parameters as outlined in Section VI-A.
 For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory.
@@
 VI-B Comparison
-The average time to solve the NLP and generate the designs presented in this section is 31.70 seconds, with a maximum of 178.41 seconds for 3mm with 3 SLRs.
+The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
 RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [60], AutoDSE [67], ScaleHLS [79] (S-HLS), and Allo [14].
@@
 PI (gmean) 2.03x 48.03x 4.92x 25.82
-TABLE II: Comparison of the throughput (GF/s) with RTL Simulation
+TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
 Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar.
@@
 $Si$ represent the statement in position $i$ in the code.
-The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
+The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
 Permutations are evident in the implementations of 3mm, Atax, and Bicg.
@@
 DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention.
-Stream-HLS [9] automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory.
-Additionally, it offers very limited opportunities for parallelism.
 Flower [5] automates FPGA dataflow development but limits parallelism.
@@
 Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
-RapidStream-TAPA [27, 28] enhances the performance of dataflow designs and automates SLR placement.
-However, it requires an optimized kernel with a dataflow structure as input.
 Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [58, 59, 60], bottleneck DSE [67], and GNN-based latency and resource estimation [66, 64, 65, 8, 68, 73, 74].
@@
 In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework.
-By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to minimize latency within stringent resource constraints.
+By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints.
 Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.

### v4

@@
 Holistic Optimization Framework for FPGA Accelerators
-Abstract
-Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization.
-Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential.
-High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives.
-However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management.
-While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
-This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space.
-Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
-I Introduction
-The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization.
-Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential.
-High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions.
-However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
-Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71].
-Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance.
-Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions.
-Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design.
-This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs).
-In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
-Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks.
-This method increases parallelism and is particularly effective for computation-bound kernels.
-Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58].
-However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained.
-While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
-Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks.
-This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out).
-Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks.
-Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
-Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use.
-A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
-Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus.
-This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation.
-The main contributions are:
-  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture.
-This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
-  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
-  • Automating our framework that is end-to-end comprehensive.
-It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
-  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14].
-Importantly, it achieves these results without requiring DSE for HLS design generation.
-II Background and Motivation
-II-A HLS optimization
-In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role.
-The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II).
-array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling.
-Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18].
-On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions.
-For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers.
-The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
-Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation.
-Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
-II-B Limitation of existing methods
-To demonstrate the limitations of the existing framework and methodology we use 3mm.
-As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
-Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication.
-AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
-Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently.
-Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved.
-Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
-Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored.
-Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
+Abstract.
+Customized accelerators have revolutionized modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization.
+Field-Programmable Gate Arrays (FPGAs) play a crucial role in this paradigm, offering unparalleled flexibility and high-performance potential.
+High-Level Synthesis (HLS) and source-to-source compilers have simplified FPGA development by translating high-level programming languages into hardware descriptions enriched with directives.
+However, achieving high Quality of Results (QoR) remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
+This paper presents Prometheus, a holistic optimization framework that integrates key optimizations—including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution—into a unified design space.
+By leveraging Non-Linear Programming (NLP) methodologies, Prometheus explores the optimization space under strict resource constraints, enabling automatic bitstream generation.
+Unlike existing frameworks, Prometheus considers interdependent transformations and dynamically balances computation and memory access.
+We evaluate Prometheus across multiple benchmarks, demonstrating its ability to maximize parallelism, minimize execution stalls, and optimize data movement.
+The results showcase its superior performance compared to state-of-the-art FPGA optimization frameworks, highlighting its effectiveness in delivering high QoR while reducing manual tuning efforts.
+High-Level Synthesis, NonLinear Programming, Compiler
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
+1. Introduction
+The rise of customized accelerators has revolutionized modern computing, driving significant improvements in energy efficiency and performance through hardware specialization.
+Among these accelerators, Field-Programmable Gate Arrays (FPGAs) have gained widespread adoption due to their flexibility, high performance, and adaptability across diverse domains, including machine learning, scientific computing, financial modeling, and embedded systems.
+Unlike fixed-function hardware such as Application-Specific Integrated Circuits (ASICs), FPGAs offer reconfigurability, enabling hardware adaptation to workload-specific requirements.
+However, achieving a good Quality of Result (QoR) on FPGAs remains a complex and resource-intensive process, requiring careful tuning of computation, memory access, and parallelism strategies.
+To address this complexity, High-Level Synthesis (HLS) tools and source-to-source compilers (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025) have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python.
+These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution.
+Despite their advantages, HLS-generated designs still require substantial manual tuning to achieve high QoR, as the compiler’s ability to optimize performance remains highly dependent on the structure and directives applied to the input code.
+1.1. Challenges in FPGA Optimization
+Despite significant advancements in HLS automation and FPGA design methodologies, achieving high-performance and resource-efficient FPGA implementations remains a challenging problem.
+This complexity arises from the intricate interactions between computation, memory access, and parallel execution strategies.
+Several key challenges must be addressed to unlock the full potential of FPGA acceleration.
+Code Transformation and Design Space Exploration: FPGA performance is highly dependent on how computations are structured and scheduled.
+Optimizations such as loop tiling, loop permutation, and loop unrolling play a crucial role in exposing parallelism and improving memory access efficiency.
+Task Concurrency and Resource Management: Maximizing throughput in FPGA designs requires efficient task concurrency, where multiple computation kernels execute in parallel.
+However, coordinating parallel tasks introduces challenges in managing resource contention, memory bandwidth, and routing complexity.
+Overlapping computations effectively demands intelligent scheduling strategies that minimize stall cycles while avoiding bottlenecks caused by limited on-chip storage and off-chip memory latency.
+Computation-Communication Overlap: FPGA acceleration is often hindered by inefficient data movement between off-chip memory (DDR/HBM) and on-chip compute units.
+Many workloads experience severe performance degradation due to memory access bottlenecks rather than computational limitations.
+Dataflow architectures aim to address this by implementing FIFO-based streaming to pipeline data transfers, but they often fail to exploit intra-task parallelism efficiently.
+In contrast, shared buffer models (Pouget et al., 2025b; Ye et al., 2022) improve data reuse, but their rigid memory management strategies make it difficult to overlap computation and communication dynamically.
+A more adaptive approach is needed to orchestrate data movement in parallel with execution, reducing overall memory stalls.
+Scalability and Multi-SLR Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability.
+However, most frameworks, including Merlin-based tools (e.g., AutoDSE, HARP, NLP-DSE) and Sisyphus, restrict designs to a single SLR, leading to underutilization of FPGA resources.
+While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity.
+Without an SLR-aware optimization strategy, large-scale tasks face a high probability of bitstream generation failures or degraded performance due to excessive inter-SLR communication overhead.
+Comprehensive Design Space Exploration: Current design space exploration (DSE) methodologies are often narrow in scope, considering only a limited subset of optimizations at a time.
+Each transformation—whether it involves loop reordering, tiling, pipelining, or data partitioning—has a profound impact on the overall performance and resource usage.
+Frameworks that optimize only one aspect of the design fail to account for the interdependencies between different optimizations, leading to suboptimal results.
+To achieve maximum efficiency, FPGA designs require a global optimization approach that integrates all possible transformations while considering the constraints imposed by memory hierarchy, computation parallelism, and hardware utilization.
+Addressing these challenges requires a unified optimization strategy that seamlessly integrates code transformation, task concurrency management, computation-communication overlap, and hardware-aware scheduling into a single framework.
+The proposed work aims to develop such an approach, ensuring high-performance FPGA implementations while maintaining synthesis feasibility.
+1.2. Contributions and Prometheus Method
+This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology.
+Unlike existing tools that optimize isolated aspects of FPGA acceleration, Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling.
+It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
+To explore the vast design space efficiently, Prometheus formulates the optimization process as a Non-Linear Programming (NLP) problem, enabling automatic selection of loop transformations, tiling strategies, pragma configurations, and memory partitioning policies.
+Unlike heuristic-based approaches, this formulation captures complex optimization interdependencies and ensures globally efficient designs.
+Additionally, Prometheus incorporates SLR-aware task scheduling, which partitions tasks across multiple Super Logic Regions (SLRs).
+This addresses a major limitation in prior works that restrict execution to a single SLR, leading to underutilization of FPGA resources.
+By integrating SLR-aware optimizations, Prometheus reduces routing congestion, improves scalability, and ensures successful bitstream generation.
+The methodology follows a structured process: First, Prometheus performs affine code analysis to identify parallelism opportunities and dependencies.
+It then generates a dataflow graph and fuses tasks that share the same outputs to minimize unnecessary communication overhead.
+Next, Prometheus constructs a comprehensive design space that includes all key optimizations outlined in Section 2, such as task concurrency via dataflow, computation-communication overlap, and adaptive parallelism, while ensuring resource constraints are met for each SLR.
+The NLP-based design space exploration model, as described in Section 4, is then employed to identify the theoretically optimal set of transformations.
+Once the best configuration is identified, Prometheus automatically generates the optimized design.
+Finally, the system produces an OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
+The key contributions of this work are:
+  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurency and computation-communication overlap while considering interdependencies between computation and data movement.
+  • Hybrid execution model that dynamically selects between shared buffering and dataflow streaming to maximize parallelism and efficient memory access.
+  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies for globally theoretical optimal performance.
+  • SLR-aware scheduling and multi-SLR partitioning to balance routing complexity and FPGA resource utilization, overcoming single-SLR limitations in prior works.
+  • End-to-end compilation and automation that generates optimized HLS-C++ code, OpenCL host code, and FPGA bitstreams with minimal manual intervention.
+  • Comprehensive performance evaluation, demonstrating superior QoR compared to AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), ScaleHLS (Ye et al., 2022), Stream-HLS (Basalama and Cong, 2025) and Allo (Chen et al., 2024).
+2. Background and Motivation
+Optimizing latency and performance in FPGA designs relies on a combination of code transformations and strategic insertion of directives within High-Level Synthesis (HLS) (Cong et al., 2018).
+This section explores various HLS optimization techniques, their role in enhancing parallelism and resource utilization, as well as their limitations and challenges.
+Table 1 presents a comparative analysis of various frameworks built on top of AMD Vitis HLS.
+These frameworks are categorized based on their underlying methodologies, including Model-Free, AI-Based, Cost Model Communication, Cost Model Computation, and NLP-Based approaches.
+The table evaluates their capabilities across key optimization techniques and objectives.
+Model Free AI Model Cost Model Communication Cost Model Computation NLP-Based
+AutoDSE (Sohrabizadeh et al., 2021) HARP (Sohrabizadeh et al., 2023) PolyOpt-HLS (Pouchet et al., 2013a) ScaleHLS (Ye et al., 2022) / POM (Zhang et al., 2024) HeteroCL (Lai et al., 2019) / Allo (Chen et al., 2024) StreamHLS (Basalama and Cong, 2025) NLP-DSE (Pouget et al., 2025a) Sisyphus (Pouget et al., 2025b) Prometheus
+Pragma Insertion Limit.
+Code Transformation (Tiling) Limit Limit
+Code Transformation (Loop Permutation) Limit Limit
+Code transformation + pragma insertion (unified)
+Task Concurrency Limit.
+Dataflow
+Computation-Communication Overlap Limit.
+Data Packing
+Padding (Communication)
+Padding (Computation)
+Hardware Aware Limit.
+Hardware-feasible Limit. Limit. Limit. Limit.
+Management of Off-Chip and On-Chip Memory Transfers
+Objective Comm + Comp Comm + Comp Comm Comp - Comp Comm + Comp Comm + Comp Comm + Comp
+Enumeration (AI, heuristics, …) Manual
+Table 1.
+Comparison of HLS-Based FPGA Optimization Frameworks.
+This table summarizes key features supported by various frameworks, including pragma insertion, code transformations, task concurrency, dataflow modeling, and hardware-awareness, categorized by their underlying optimization strategies.
+2.1. HLS Optimization
+2.1.1. Pragma Insertion
+HLS optimizations rely extensively on a range of directives that control loop execution, data access patterns, and computation scheduling to maximize performance.
+Among these, three fundamental pragmas—unroll, pipeline, and array partitioning—are particularly critical for improving parallelism and reducing execution latency.
+The unroll directive enables the expansion of loop iterations in the hardware design, allowing multiple iterations to execute in parallel.
+By eliminating loop control overhead and enabling greater resource utilization, unrolling can significantly boost throughput.
+However, its effectiveness is limited by FPGA resource constraints, such as available DSP slices, BRAM, and routing congestion.
+The pipeline pragma restructures loop execution to allow overlapping operations, thereby reducing the initiation interval (II)—the number of cycles required to launch consecutive iterations.
+This directive is essential for achieving high throughput in applications with iterative computations, such as matrix multiplications and stencil operations.
+Careful tuning of pipeline depth and initiation interval is necessary to balance resource usage while avoiding performance bottlenecks due to memory access contention.
+Array partitioning is another crucial optimization that enhances memory access parallelism.
+By splitting large arrays into smaller banks stored in separate on-chip BRAM blocks or distributed memory structures, this pragma allows simultaneous data accesses, facilitating more effective unrolling and pipelining.
+Without partitioning, memory conflicts and bandwidth limitations can restrict performance, especially with multiple concurrent memory accesses.
+Beyond these fundamental pragmas, additional directives such as loop flattening, and dependency pragmas can further optimize execution.
+Loop flattening combines nested loops into a single loop to improve hardware efficiency.
+Dependency pragmas help manage data dependencies to ensure efficient scheduling without unnecessary synchronization delays.
+Pragma insertion has been extensively studied using various methodologies to optimize performance and resource utilization in HLS designs.
+These approaches include bottleneck analysis, as employed by AutoDSE (Sohrabizadeh et al., 2021), performance estimation through Graph Neural Networks (GNN) in frameworks like Harp (Sohrabizadeh et al., 2023), and cost model optimization leveraging Non-Linear Programming (NLP) in NLP-DSE (Pouget et al., 2025a, 2024).
+Each method presents distinct advantages and limitations.
+AutoDSE achieves full accuracy by running the HLS compiler for each configuration, ensuring precise performance measurements.
+However, this approach is computationally expensive and time-consuming, making it impractical for rapid design exploration.
+HARP, on the other hand, can estimate performance across configurations without direct compilation, significantly reducing evaluation time.
+However, it requires extensive training datasets and fine-tuning for each new kernel to accurately model compiler behavior.
+While this approach enables learning-based adaptation to different designs, its effectiveness depends heavily on data quality and model calibration.
+Conversely, NLP-DSE explores the entire design space efficiently by formulating the optimization as a Non-Linear Programming problem.
+This method allows for rapid exploration and selection of theoretical optimal configurations.
+However, it relies on a theoretical cost model, which may introduce inaccuracies if the compiler behaves unpredictably or if certain optimizations are not accurately captured by the model.
+Consequently, while NLP-DSE provides a fast and scalable solution, its reliability depends on the fidelity of the cost model to actual HLS compilation behavior.
+While pragma insertion is a powerful optimization technique, its effectiveness is inherently dependent on the interaction between memory hierarchy, computational resources, and FPGA-specific constraints.
+Moreover, its impact is tightly coupled with the program’s execution schedule, meaning that without appropriate code transformations, the benefits of pragma insertion remain limited.
+Effective optimization requires a synergy between pragma directives and structural modifications to the code to fully exploit FPGA parallelism and resource utilization.
+2.1.2. Code Transformation
+Code transformations are fundamental for optimizing execution on FPGAs.
+While pragma insertion plays a crucial role in enhancing performance, it must be complemented by effective code transformations to fully exploit FPGA resources.
+Techniques such as loop reordering (permutation), task fusion, and data tiling enhance data locality and increase parallelism, thereby improving both computational efficiency and memory access patterns.
+Various code transformation strategies have been developed specifically for FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018).
+While transformations originally designed for CPUs and GPUs achieve substantial performance gains by optimizing for their respective architectures, they do not inherently align with FPGA requirements, which prioritize fine-grained parallelism and efficient resource utilization.
+Several studies have utilized Pluto (Bondhugula et al., 2008), a leading compiler framework originally designed for CPU optimizations, to transform FPGA kernels (Zhao et al., 2022; Zhao and Cheng, 2021).
+While Pluto excels at tiling and minimizing dependencies to enhance memory reuse, its direct application to FPGA optimization is constrained due to the fundamental differences in optimization strategies required for CPUs and FPGAs.
+Unlike CPUs, where memory hierarchy and cache locality are primary concerns, FPGA optimization focuses on maximizing parallelism, minimizing resource contention, and efficiently utilizing on-chip memory, making Pluto’s conventional transformation techniques less effective in this context.
+Conversely, studies such as (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on code transformations tailored to specific FPGA performance goals.
+The work in (Pouchet et al., 2013b) aims to reduce communication overhead between off-chip and on-chip memory, achieving superior Quality of Results (QoR) for memory-bound kernels.
+Meanwhile, research efforts in (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on optimizing pipelining strategies to maximize instruction-level parallelism and resource utilization.
+Sisyphus (Pouget et al., 2025b) introduces a unified approach by integrating code transformation and pragma insertion into a single optimization problem.
+By formulating this as a Non-Linear Programming (NLP) problem, Sisyphus efficiently explores the design space to identify theoretical optimal configurations, streamlining FPGA acceleration while maintaining a balance between computation and memory access efficiency.
+2.1.3. Task Concurrency
+HLS tools like Vitis HLS support the dataflow pragma, which structures computations into actors that communicate through FIFO queues.
+This approach allows overlapping execution of multiple tasks, significantly reducing overall latency .
+By enforcing a producer-consumer model, dataflow scheduling enables each task to process data as soon as it becomes available, rather than waiting for an entire stage to complete, which is particularly beneficial.
+For computational kernels such as 3mm (Listing 3), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced.
+This overlapping of execution helps maximize throughput and minimize idle time for computing units.
+Additionally, dataflow optimizations facilitate task-level parallelism, allowing independent tasks to run concurrently across multiple compute resources, such as DSP blocks and BRAM, ensuring efficient utilization of available FPGA resources.
+However, despite its advantages, pure dataflow parallelism presents several challenges.
+One major limitation is intra-task parallelism, which is constrained by the reliance on FIFO-based communication.
+Since each FIFO can only transfer up to 512 bits per cycle (as this is the maximum off-chip memory bitwidth supported), this inherently restricts the amount of data that can be processed concurrently within a task.
+To further enhance intra-task parallelism, alternative approaches must be explored.
+Stream-HLS (Basalama and Cong, 2025) attempts to address this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip.
+While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory.
+By using $n$ FIFOs to transfer $n$ data elements in parallel between two tasks, the method significantly increases resource consumption without proportionally improving efficiency.
+The additional FIFOs complicate the design, potentially leading to routing congestion and excessive hardware overhead.
+Moreover, modern FPGAs feature up to 32 off-chip memory banks, making the multi-FIFO approach impractical for handling off-chip data transfers.
+A more effective and scalable strategy must be developed to optimize intra-task parallelism while ensuring efficient communication between off-chip and on-chip memory, without introducing unnecessary complexity or resource constraints.
+2.1.4. Shared Buffering
+Shared buffering is a critical technique in FPGA memory optimization, enabling efficient data access and reuse across multiple computational units.
+It can be employed at a global level, as seen in frameworks like AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), and ScaleHLS (Ye et al., 2022), or within individual dataflow tasks to enhance execution efficiency.
+This approach involves preloading data buffers into on-chip memory, allowing multiple computations to access shared data without redundant transfers.
+By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance.
+It is particularly beneficial in applications which are computation bound.
+However, shared buffering presents challenges in maintaining concurrency.
+While it enables efficient data reuse, it can introduce synchronization overhead, especially when multiple tasks attempt to access the same memory region simultaneously.
+Managing concurrent access requires arbitration mechanisms, which can lead to increased latency and potential bottlenecks.
+Additionally, routing congestion can occur due to high interconnect demands, limiting scalability in complex FPGA designs.
+One limitation of shared buffering is its impact on initiation interval (II) in pipelined architectures.
+Unlike dataflow-based designs that rely on FIFO queues for seamless data streaming, shared buffering requires explicit read and write coordination, which may introduce stalls if not carefully managed.
+Moreover, FPGA resource constraints, such as limited BRAM and URAM availability, impose restrictions on buffer size and allocation strategies.
+To address these challenges, advanced techniques such as double-buffering, memory partitioning, and adaptive scheduling have been explored.
+Double-buffering enables overlapping computation with data transfer, reducing idle cycles.
+Memory partitioning distributes data across multiple banks to alleviate contention, while adaptive scheduling dynamically assigns buffer access based on task priority and workload demands.
+By effectively integrating shared buffering with intelligent memory management strategies, FPGA designs can achieve a balance between computational parallelism and efficient memory access.
+Future advancements should focus on automated buffer allocation techniques and dynamic access pattern optimization to further enhance performance in high-level synthesis (HLS) workflows.
+2.1.5. Computation-Communication Overlap
+Overlapping computation and communication is crucial for high-performance FPGA designs.
+Techniques such as double buffering (ping-pong buffering) and advanced data tiling help mask communication latency while keeping computational units busy.
+Managing data transfers efficiently between on-chip and off-chip memory ensures that processing units remain active without waiting for data.
+2.1.6. Data Packing and Padding
+Modern FPGA architectures support high-bandwidth data transfers (up to 512-bit wide on AMD/Xilinx FPGAs).
+Data packing and padding optimize memory alignment, reducing the number of required memory cycles.
+For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles, making efficient packing crucial for minimizing overhead.
+To fully exploit data packing, padding must be considered to enable even faster data transfers.
+Additionally, padding is valuable for achieving finer control over parallelism and resource utilization by expanding the available design space for loop unrolling, thereby enhancing computational throughput and efficiency.
+Padding for Communication
+Padding must be considered to increase the flexibility of data packing.
+The original size of the data may impose restrictions on packing efficiency, but by introducing padding, we create a larger space that may allow for a more efficient transfer.
+However, padding is not a free optimization, as increasing it also increases the amount of data that needs to be transferred.
+For example, in the code presented in Listings 1 and 2, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth.
+Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$.
+However, by introducing padding with $P=2$ and adjusting $J$ to $192$ ($190+2$), the data alignment enables a significantly higher transfer rate of $512$ bits per cycle, as $192\times 32$ is now perfectly divisible by $512$.
+This optimization maximizes memory bandwidth utilization, reducing transfer latency and improving overall data throughput.
+Padding for Computation
+Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 1 to Listing 2.
+We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra ressource to only compute a partial tile which would correspond to the execution of the last tile which is not complete.
+For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
+But we restrict the unroll factor to be divisible of the trip count we restrict considerably the space of possibilities.
+So the solution is to pad which allow to considerably increase the space of possibilities.
+Hence with a trip count of $190$ the possible unroll factor are include in $UF\in\{1,2,5,10,19,38,95,190\}$ but with padding of 2 (a trip count of 192) the space become $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}$.
+And of course all pading can be considered.
+Padding enables finer control over unrolling and resource utilization, allowing for more flexible and efficient hardware implementation while optimizing computational parallelism.
+1for (i = 0; i < I; i++)
+2 for (j = 0; j < J; j++)
+3 for (k = 0; k < K; k++)
+4 C[i][j] += A[i][k] * B[k][j];
+Listing 1: Baseline Implementation of Matrix Multiplication in C.
+This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
+1for (i0 = 0; i0 < I0; i0++)
+2 for (j0 = 0; j0 < J0; j0++)
+3 for (k0 = 0; k0 < K0; k0++)
+4 for (i1 = 0; i1 < I1; i1++)
+5#pragma HLS unroll
+6 for (j1 = 0; j1 < J1; j1++)
+7#pragma HLS unroll
+8 for (k1 = 0; k1 < K1; k1++)
+9#pragma HLS unroll
+10 C[i0*I1 + i1][j0*J1 + j1] += A[i0*I1 + i1][k0*K1 + k1] * B[k0*K1 + k1][j0*J1 + j1];
+Listing 2: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation.
+This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
+2.2. Performance and Hardware Considerations
+2.2.1. Performance Evaluation
+Performance evaluation in HLS is a multi-stage process that assesses different aspects of a design’s efficiency and feasibility.
+It is typically conducted at three levels: estimation (e.g., Vitis HLS reports), RTL simulation, and FPGA-based evaluation.
+Each of these stages provides critical insights, but they vary in accuracy, speed, and resource requirements.
+  1. (1)
+    Estimation (HLS Reports): The first level of evaluation relies on estimation tools provided by HLS compilers such as Vitis HLS.
+These reports generate quick approximations of key performance metrics, including latency, resource utilization (DSPs, LUTs, BRAMs), and throughput.
+While these estimations are useful for early-stage design exploration, they do not account for low-level placement and routing effects, which can significantly impact real-world performance.
+The optimistic assumptions of HLS reports often fail to capture routing congestion, memory access delays, and other architectural bottlenecks.
+  2. (2)
+    RTL Simulation: To achieve more accurate performance insights, RTL simulation is performed after HLS compilation.
+This stage involves generating synthesizable RTL code, which is then simulated using hardware description language (HDL) tools.
+RTL simulation provides cycle-accurate performance metrics, allowing designers to analyze pipeline behavior, data dependencies, and execution timing.
+However, while RTL simulation models hardware more precisely than HLS reports, it does not incorporate real-world FPGA constraints such as clock tree synthesis, interconnect delays, and power distribution, which can affect the final implementation.
+  3. (3)
+    FPGA-Based Evaluation: The most accurate method of performance evaluation involves deploying the design onto an actual FPGA.
+This process includes place-and-route, bitstream generation, and execution on hardware.
+FPGA-based evaluation provides real execution metrics such as operating frequency, power consumption, and effective memory bandwidth.
+It also reveals practical challenges, including timing closure issues, placement inefficiencies, and resource contention that cannot be detected in earlier stages.
+This level of evaluation is essential for validating Quality of Results (QoR) and ensuring the design meets real-world constraints.
+The transition from RTL simulation to FPGA implementation often introduces additional challenges, such as increased routing complexity, unexpected timing violations, and suboptimal resource utilization.
+These discrepancies arise due to the abstraction gap between HLS-generated code and final FPGA hardware constraints.
+To mitigate these issues, it is essential to develop robust tools that enable rapid design regeneration by modifying specific configurations within a well-defined portion of the design space.
+Such tools would enable engineers to efficiently iterate on design parameters without requiring a complete code regeneration, which can be time-consuming and may result in a significantly different design that still fails to generate a valid bitstream.
+Having the ability to selectively modify only the congested parts of the design while preserving the rest of the configuration would be highly valuable, ensuring faster convergence toward an optimized and feasible FPGA implementation.
+2.2.2. Hardware Awareness and Resource Constraints
+Many existing studies conclude their evaluations at the Vitis HLS report or RTL simulation stage, overlooking the critical impact of placement and routing constraints.
+However, as the design progresses toward bitstream generation, the available design space becomes increasingly constrained.
+Addressing hardware limitations early in the HLS process is essential to avoid costly design iterations and ensure convergence toward a feasible and deployable FPGA implementation.
+Additionally, Super Logic Region (SLR)-aware optimizations are crucial for multi-SLR FPGA architectures.
+A Super Logic Region (SLR) is a physically distinct section of an FPGA die in Stacked Silicon Interconnect (SSI) technology.
+Each SLR contains a subset of the device’s computational resources, including DSP, LUT and FF, similar to those in monolithic FPGA devices.
+Effectively mapping tasks across SLRs is critical for balancing resource utilization and minimizing routing congestion, as inter-SLR communication introduces additional latency and can become a performance bottleneck.
+By integrating SLR-aware task partitioning, optimized scheduling, and efficient data movement strategies, FPGA designs can achieve higher scalability, improved parallelism, and enhanced synthesis feasibility.
+Ensuring that computation and memory access patterns align with the physical structure of multi-SLR FPGAs is essential for achieving high QoR while maintaining timing closure and efficient resource allocation.
+Some recent frameworks, such as RapidStream-TAPA (Guo et al., 2023a, b) and PASTA (Khatti et al., 2023), aim to improve the design scalability and performance of HLS programs on modern multi-die FPGAs through hardware-aware co-optimization of HLS and physical design.
+While RapidStream-TAPA focuses on latency-insensitive, FIFO-based communication and exploits coarse-grained floorplanning and pipelining for timing closure and parallel compilation, PASTA extends this model to include buffer-based inter-task communication using a generalized channel abstraction.
+Despite their advances, both frameworks still require a manually optimized task-parallel input code, which demands substantial expertise in HLS and hardware design, limiting their accessibility for non-expert users.
+2.3. Challenges
+While existing HLS optimizations provide substantial performance improvements, several challenges remain.
+One of the primary limitations is that many prior works restrict their optimization space to specific techniques or separate their optimization processes into multiple independent steps.
+This fragmented approach can lead to incoherent design decisions, where optimizations applied at one stage may not align with or may even counteract those applied at subsequent stages.
+Refer to Table 1 for a detailed comparison of these approaches.
+A common oversimplification in many frameworks is the assumption that all data resides on-chip.
+While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware.
+This assumption prevents frameworks from incorporating crucial tiling strategies, which are essential for optimizing data locality, reducing memory access overhead, and ensuring efficient off-chip communication.
+When off-chip memory management is treated as a separate optimization step, decisions made in earlier phases may restrict the effectiveness of later memory optimizations, leading to suboptimal performance and resource utilization.
+The limitations of current HLS optimization frameworks that rely on shared buffering are exemplified through the 3mm kernel.
+It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication.
+As illustrated in Listing 3, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap.
+However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
+Another major challenge is balancing parallelism and routing complexity.
+Excessive parallelization can lead to routing congestion, increased place-and-route time, and ultimately, failure to generate a valid bitstream.
+Current methodologies lack adaptive mechanisms to dynamically adjust parallelism levels based on available routing resources and FPGA architecture constraints.
+Without such mechanisms, achieving high-performance designs often requires manual intervention and iterative fine-tuning.
+While pragma-based optimizations such as unroll and pipeline significantly improve performance, their effectiveness is inherently tied to the underlying code structure.
+Without proper transformations to expose parallelism and optimize memory access patterns, pragma insertion alone cannot fully unlock the potential of FPGA acceleration.
+Addressing these challenges requires a holistic approach that integrates memory-aware optimizations, automated design space exploration, and adaptive scheduling techniques.
+Future research should focus on frameworks that jointly optimize data placement, computational parallelism, and resource utilization while considering real hardware constraints.
+By developing more cohesive and intelligent HLS optimization strategies, FPGA-based acceleration can achieve greater efficiency, scalability, and deployment feasibility.
+Model-free approaches, such as AutoDSE (Sohrabizadeh et al., 2021) and AI-based frameworks like HARP (Sohrabizadeh et al., 2023; Wu et al., 2023, 2022a), rely on either HLS compilers or AI models to predict performance and resource utilization.
+However, these methods are currently limited to pragma insertion and space enumeration.
+Even though AI-based models provide fast estimations (on the order of milliseconds), exhaustively exploring large design spaces remains infeasible.
+Our previous work, NLP-DSE (Pouget et al., 2025a), addresses this limitation by leveraging NLP techniques to efficiently explore pragma configurations.
+However, like the other approaches, it remains restricted to pragma insertion.
+PolyOptHLS (Pouchet et al., 2013a) focuses on optimizing designs by minimizing memory transfers.
+While reducing communication latency can significantly improve performance, it does not always lead to optimal results.
+In many cases, the primary bottleneck may shift from communication to computation, limiting overall efficiency and potentially degrading the quality of results (QoR).
+This approach, therefore, lacks a balanced optimization strategy that considers both computation and communication trade-offs.
+ScaleHLS (Ye et al., 2022) and POM (Zhang et al., 2024) extend optimization beyond pragma insertion by incorporating code transformations.
+However, their exploration space is constrained by the assumption that data is already on-chip.
+Additionally, their transformations are based on heuristics, such as permuting the reduction loop to the outermost level.
+While they employ a cost model to minimize computation latency, they still rely on exhaustive enumeration of possible configurations.
+Allo (Chen et al., 2024) aims to reduce development time and enhance design quality.
+However, it still requires significant manual intervention from an expert, limiting its usability for non-specialists.
+Our previous work, Sisyphus (Pouget et al., 2025b), enables both code transformations and pragma insertion but is restricted to optimizing a single task.
+It lacks support for optimizations that dataflow techniques can provide, which would be necessary for broader efficiency improvements.
+Stream-HLS (Basalama and Cong, 2025) effectively leverages dataflow pragmas by selecting a good loop ordering strategy to maximize streaming efficiency.
+However, it imposes constraints on the design space by assuming that data is on-chip.
+Moreover, its approach to increasing parallelism relies on multiple FIFOs, which is not generalizable when off-chip memory access is required.
+A significant limitation of all these frameworks is that none of them account for hardware constraints in their exploration space.
+While this simplifies the exploration process, it reduces real-world applicability.
+Additionally, frameworks such as Merlin-based tools (AutoDSE, HARP, NLP-DSE) and Sisyphus can generate all necessary components, including off-chip memory, for bitstream generation.
+However, they are constrained to a single Super Logic Region (SLR), leading to under-utilization of the FPGA board.
+Although generating a design that spans multiple SLRs is theoretically possible, it significantly increases the risk of bitstream generation failures.
+Even if the bitstream is successfully generated, timing violations often degrade performance.
 1for (i = 0; i < 180; i++) // MM 1
@@
 15 G[i][j] += E[i][k] * F[k][j];} // S5
-Listing 1: Code of 3mm
-Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs.
-For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers.
-However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
-As 3mm is computation-bound, increasing parallelism is vital.
-However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation.
-Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
-Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR.
-Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
-Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency.
-These gaps can lead to suboptimal performance, particularly in data-intensive scenarios.
-Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors.
-However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops.
-Expanding unroll factor options could better balance parallelism and performance.
-For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74).
-These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method.
-However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
-II-C Overview of Prometheus
-We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently.
-Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
-The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel.
-The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs.
-Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1.
-We can see an example on Listing 3.
-Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded.
-Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
-It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
+Listing 3: Reference Implementation of the 3mm Kernel from PolyBench.
+This listing shows the baseline version of the 3mm benchmark, comprising three matrix multiplication operations.
+The first two multiplications are independent and produce intermediate results that serve as inputs to the final matrix multiplication.
+2.4. Overview of Prometheus
+We introduce Prometheus, a unified optimization framework designed to efficiently navigate complex design spaces using a Non-Linear Programming (NLP) cost model.
+Prometheus streamlines FPGA design development by providing an intuitive, engineer-friendly interface with tunable parameters, including tile size, array partitioning, and loop unrolling factors.
+The framework integrates all the optimization techniques discussed in previous sections, ensuring a comprehensive and cohesive approach to performance enhancement.
+Additionally, Prometheus is SLR-aware, enabling intelligent task distribution across Super Logic Regions (SLRs).
+This capability not only simplifies bitstream generation but also optimizes the trade-off between computational performance and resource utilization.
+Prometheus C++ Code Intermediate Representation (dependencies, trip counts) Task-flow Graph
+1234
+Formulate a Non-Linear Problem encoding the full optimization design space All Parameters of the Spacee.g.
+data-tile of array $A$ is $\texttt{32}\times\texttt{64}$Files for HLS CompilerHLS-C++, OpenCL, Config Bitstream with HLS Compiler
+Extract with PoCC GenerateSolve NLP Generate
+Figure 1.
+Overview of the Prometheus Framework Workflow.
+This diagram illustrates the end-to-end pipeline of the Prometheus optimization framework, starting from C++ source code and proceeding through intermediate representation extraction, task-flow graph generation, Non-Linear Programming (NLP)-based design space exploration, and final compilation into FPGA bitstreams using HLS compilers.
+Group Parameter Description
+Design Variables
+Memory Bit Width ($BW_{a}$) Width of memory transfer in bits;
+Data-tile Size (modeled via Transfer Level $t_{a,l}$ in Section 4) Size of the data tile of array $a$ transferred either from off-chip to on-chip memory or between tasks.
+The transfer level determines this size.
+Data-tile Reuse Size (modeled via Reuse Level $d_{a,l}$ in Section 4) Size of the data tile of array $a$ reused within a task. The reuse level dictates this tile size.
+Buffering ($N_{a}$) Number of buffers (2 for double buffering, 3 if read & write), enabling overlap of load/compute/store.
+Communication Padding Additional data added to enable wider burst transfers and better bandwidth utilization.
+Parallelism Unroll Factors ($TC^{l}_{\text{intra}}$) Unroll factor for each task, determined through tiling, where the intra-tile loop is fully unrolled to expose fine-grained parallelism.
+Array Partitioning ($AP_{a,d}$) Number of partitions for array $a$ along dimension $d$ to support parallel accesses.
+Compute Padding Padding applied to loop trip counts to increase legal unroll factors.
+Code Structure Loop Permutations Legal reordering of loops to optimize data locality and parallelism; synchronized across fused statements.
+Tiling Splitting loops into inter- and intra-tile levels for optimization and pipelining.
+SLR Assignment ($slr_{t}$) Mapping of each task $t$ to a Super Logic Region (SLR) to enable spatial task distribution.
+Design Constraints
+Resource Limits Max Array Partitioning Total partitions across all arrays must not exceed architectural limits.
+DSP Budget Maximum DSP slices per SLR; used to limit total DSP usage by pipelined/unrolled tasks.
+On-Chip Memory Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
+Table 2.
+Design Variables and Architectural Constraints in the Prometheus Optimization Space.
+This table outlines the key configurable parameters used in Prometheus—including loop tiling, unrolling, array partitioning, and buffer management—along with hardware constraints such as resource limits and memory capacity.
+The cost model also accounts for concurrent task execution enabled by the dataflow pragma, as well as computation-communication overlap achieved through automatic double or triple buffering.
+These elements jointly define the valid and efficient design space explored by Prometheus.
+Figure 1 illustrates the workflow of Prometheus.
+The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information.
+This extraction is performed using PoCC (Pouchet, 2023), which provides all necessary details for analysis.
+Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem.
+Solving this NLP problem determines the theoretical optimal parameters for the design space.
+Once the NLP solver provides these parameters, we have all the required information to proceed.
+At this stage, we automatically generate the HLS-C++ code along with all the necessary files to produce the FPGA bitstream, ensuring an efficient and fully automated compilation process.
+To support efficient exploration of the design space, Prometheus defines a set of tunable parameters and architectural constraints that capture the key aspects of FPGA acceleration.
+Table 2 provides a detailed overview of these parameters, including memory-related configurations (e.g., bitwidth, tiling, and reuse levels), parallelism controls (e.g., loop unrolling and array partitioning), and structural transformations (e.g., loop permutations and SLR assignments).
+In addition, the table outlines hardware constraints such as DSP budgets, on-chip memory limits, and maximum legal array partitioning.
+These variables and constraints jointly define the feasible design space that the NLP-based optimization engine systematically explores to generate high-performance, hardware-aware FPGA implementations.
+The pseudo-code in Listing 4 demonstrates how Prometheus processes the 3mm kernel (Listing 3).
+Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
+Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 3. An example of such a task is illustrated in Listing 5.
+Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
+Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 3), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
+The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
 1/*************** Fused Task 0 ****************/
@@
 6 float E[10][32];
-7 task_0(E); // S0 + intra-tile loops
+7 task0(E); // S0 + intra-tile loops
 8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
 9#pragma HLS pipeline II=3
-10 task_1(E, A, B); // S1 + intra-tile loops
+10 task1(E, A, B); // S1 + intra-tile loops
 11 store_E(E); sent_E(E);
@@
 17 float F[19][32];
-18 task_2(F); // S2 + intra-tile loops
+18 task2(F); // S2 + intra-tile loops
 19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
 20#pragma HLS pipeline II=3
-21 task_3(F, C, D); // S3 + intra-tile loops
+21 task3(F, C, D); // S3 + intra-tile loops
 22 store_F(F); sent_F(F);
@@
 30 receive_E(E);
-31 task_4(G); // S4 + intra-tile loops
+31 task4(G); // S4 + intra-tile loops
 32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
 33#pragma HLS pipeline II=3
-34 task_5(G, E, F); // S5 + intra-tile loops
+34 task5(G, E, F); // S5 + intra-tile loops
 35 store_G(G);
-Listing 2: Pseudo Code of 3mm generated by Prometheus
-Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks.
-Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready.
-These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels.
-Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
-III Problem Formulation
+Listing 4: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus.
+This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.
+Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance.
+Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
+To assess the effectiveness of our proposed framework, we evaluate the performance of the 3mm kernel using several state-of-the-art HLS-based FPGA optimization tools.
+As shown in Table 3, Prometheus significantly outperforms existing frameworks in terms of throughput, achieving 368.36 GF/s.
+This represents a substantial improvement over Sisyphus (178.97 GF/s) and Stream-HLS (174.00 GF/s), and far exceeds the results of Allo, ScaleHLS, and AutoDSE.
+The superior performance of Prometheus stems from its holistic design space exploration strategy, which integrates loop transformations, memory tiling, concurrent task execution, and SLR-aware scheduling to deliver optimized and hardware-feasible designs.
+Metric Prometheus Sisyphus Stream-HLS Allo ScaleHLS AutoDSE
+Throughput (GF/s) 368.36 178.97 174.00 60.40 43.04 1.74
+Table 3.
+Measured Throughput of the 3mm Kernel Using Various FPGA Frameworks.
+This table presents the runtime throughput (in GF/s) obtained from RTL simulation of the 3mm kernel across multiple frameworks.
+Prometheus is shown to outperform other methods, including Sisyphus, Stream-HLS, and AutoDSE.
+3. Code Transformation
 In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time.
@@
 While this constraint limits graph configurations and may increase resource usage, it reduces overall latency.
-To support this structure, we inline [15] each function in the input code to generate the required acyclic graph.
+To support this structure, we inline (Chen et al., 1993) each function in the input code to generate the required acyclic graph.
 Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently.
-To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
-Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
-ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved.
-After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies.
+To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section 4.
+3.1. Dependency Graph Creation
+Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop.
+ISCC (Verdoolaege, 2011) verifies the legality of these transformations, ensuring dependencies are preserved.
+After achieving full distribution, we construct a dependency graph using PoCC (PoCC, [n.
+d.]).
+PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph.
+In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies.
 Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties.
+This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once.
 If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
 Prometheus’s solution space includes:
-Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes.
+3.2. Data-tiling and Padding
+In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes.
 Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets.
 Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied.
-If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop.
+For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
 To optimize memory transfers, padding is applied to arrays in two ways.
@@
 Tile sizes are consistent within a fused task but vary between tasks.
-For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
-Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop.
-Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner).
-If permutation is infeasible, the inter-tile loop remains in its legal position.
-In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism.
+In Listing 4, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
+3.3. Fine-grained Parallelism
+Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops.
+Using ISCC (Verdoolaege, 2011), we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner).
+If permutation is not feasible, the inter-tile loop retains its legal position.
+For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction.
+For instance, in Listing 4, the inter-tile loops on lines 3 and 5 iterate over task0 and task1.
+The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip.
 Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access.
 Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency.
-For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$.
-Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once.
-Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data.
-Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
-Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile.
+For instance, in Listing 4, additions take 3 cycles, resulting in $II=3$.
+3.4. Loop Order
+Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile.
 We place the inter-tile reduction loops directly above the task and are pipelined.
@@
 This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops.
-This order will subsequently be determined by the NLP, as described in Section IV.
-As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
-Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers.
+This order will subsequently be determined by the NLP (cf.
+Section 4).
+As shown in Listing 4, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 3.
+3.5. Automatic Overlapping of Communication and Computation
+To overlap communication and computation, we use on-chip buffers.
 The buffer size and data transfer location are determined by various options explored via NLP (cf.
-Section IV).
+Section 4).
 Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start.
@@
 If defined or transferred before loops, $l=0$.
-Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written.
-Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
-Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks  Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution.
+Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written.
+Following (Cong et al., 2018), we perform an initial load, then overlap loading the next tile with computing the current one.
+3.6. Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks
+Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution.
 This concurrency significantly increases overall performance.
-Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding.
+3.7. Memory Transfer
+Communication latency is reduced by transferring more data per cycle using increased bit width with padding.
 Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic.
-For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
-IV NLP Formulation
+For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
+4. NLP Formulation
 To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency.
-This approach builds upon the methodology proposed by Pouget et al.
-[59, 58, 60], which we have adapted to address our specific requirements and constraints.
-In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section III.
-We employ PoCC [54] to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement.
-ISCC [69] then generates all legal permutations for each loop body.
-Constant Description
+This approach builds upon the methodology proposed in our previous work (Pouget et al., 2025a, 2024, b), which we have further adapted to meet the specific requirements and constraints of our current framework.
+In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section 3.
+We employ PoCC (PoCC, [n.
+d.]) to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement.
+ISCC (Verdoolaege, 2011) then generates all legal permutations for each loop body.
+Table 4 delineates the sets, variables, and constants utilized in our NLP formulation.
+Constants Description
 $II_{l}$ II of the loop $l$
@@
 $SLR$ Number of SLR available for the FPGA used
-Variable Description
+Variables Description
 $TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
@@
 $t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
-$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
+$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
 $slr_{t}$ ID of the SLR use by the task $t$
-Set Description
+Sets Description
 $\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
@@
 $\mathcal{T}$ The set of tasks in the dataflow
-TABLE I: Overview of the constant, variable and set employed in formulating the NLP
-IV-A Constraints
-We now describe the constraints by using the code of Listings 1, 2 and 3.
+Table 4.
+Mathematical Notation for Constants, Variables, and Sets in the NLP-Based Optimization Model.
+This table defines the formal notation used in Prometheus’ Non-Linear Programming (NLP) model for design space exploration, including loop trip counts, bitwidths, SLR mappings, resource limits, and legal transformations.
+4.1. Constraints
+We now describe the constraints by using the code of Listings 3, 4 and 5.
 1for (int j1 = 0; j1 < 32; j1++) {
@@
 8 F[i1][j1] += C[i][k] * D[k][j1];}
-Listing 3: Pseudo Code of the task_3
-Data-tiling and Unroll Factor  The intra-tile transformation, as explained in Section III, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities.
+Listing 5: Implementation of Task 3 in the 3mm Kernel (from Listing 4).
+This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism.
+The unrolling enables concurrent execution of loop iterations.
+4.1.1. Data-tiling and Unroll Factor
+The intra-tile transformation, as explained in Section 3, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities.
 Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities.
 The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver.
-For instance, in Listing 3, for the array $F$, the loop $j$ (line 7 in Listing 1) has been split into $j0$ (line 14 in Listing 2) and $j1$ (intra-tile).
+For instance, in Listing 5, for the array $F$, the loop $j$ (line 7 in Listing 3) has been split into $j0$ (line 14 in Listing 4) and $j1$ (intra-tile).
 The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
-$\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$ (1)
-$(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
-Bit Width  $\mathcal{B}$ represents the possible bit width for the data type.
+(1) $\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$
+(2) $(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$
+4.1.2. Bit Width
+$\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type.
 Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$.
@@
 Therefore, it has a bit width of 16.
-$\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
-Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement.
+(3) $\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$
+4.1.3. Permutation
 Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task.
@@
 However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
-$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
 $\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
-$\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$ (5)
-Transfer and Reuse  Equation 6 permits the selection of a single level where each array can be defined (and reused) and transferred.
+(4) $\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$
+4.1.4. Transfer and Reuse
+Equation 5 permits the selection of a single level where each array can be defined (and reused) and transferred.
 Equation 6 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
-The array $E$, defined on line 24 in Listing 2, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops).
+The array $E$, defined on line 24 in Listing 4, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops).
 However, it is transferred under the loop $i0$, so $t_{E,i0}=1$.
@@
 We cannot transfer E under loop $i0$ if E is defined under $k0$.
-Similarly, in Listing 2, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
-$\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$ (6)
-$\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$ (7)
-On-chip Memory  Equation 8 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
-$\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$ (8)
+Similarly, in Listing 4, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
+(5) $\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$
+(6) $\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$
+4.1.5. On-chip Memory
+Equation 7 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
+(7) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$
 with $N_{a}$ being the number of double buffer for the array $a$.
-Array Partitioning  Equation 9 limits the maximum partitioning of each array.
+4.1.6. Array Partitioning
+Equation 8 limits the maximum partitioning of each array.
 This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs.
 Equation 9 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
-Array D in Listing 3 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$).
+Array D in Listing 5 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$).
 Therefore, the total number of partitions needed is $3\times 32=96$.
@@
 However, this value must be less than or equal to $max_{part}$.
-$\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$ (9)
-$\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$ (10)
-DSP Utilization  Equation 11 constrains the number of DSPs used based on the available DSP resources.
-In contrast to [59, 58, 60], we utilize pessimistic DSP utilization.
+(8) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
+(9) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$
+4.1.7. DSP Utilization
+Equation 10 constrains the number of DSPs used based on the available DSP resources.
+In contrast to (Pouget et al., 2025a, 2024, b), we utilize pessimistic DSP utilization.
 This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously.
@@
 However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
-$\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$ (11)
-SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12).
-Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
-$\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (12)
-IV-B Objective Function
+(10) $\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$
+4.1.8. SLR Selection
+For each task, the NLP determines the SLR on which the task will be implemented (Equation 11).
+Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
+(11) $\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$
+4.2. Objective Function
 The objective function models the dependency graph of the fused task.
@@
 $\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
-T0T1T2T3T4T5
-Figure 1: Dataflow graph of 3mm maximally distributed
-The latency of the intra-task for a statement $s$ is determine similarly to [59, 58, 60] with Equation 13.
+PrometheusT0T1T2T3T4T5
+Figure 2. Maximally distributed dataflow graph of the 3mm kernel, where each node represents a task and each edge indicates a data dependency between tasks.
+The latency of the intra-task for a statement $s$ is determine similarly to (Pouget et al., 2025a, 2024, b) with Equation 12.
 And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
-$Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$ (13)
-$Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$ (14)
+(12) $Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$
+(13) $Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$
 For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication.
 The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
-$Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$ (15)
+(14) $Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$
 And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
-V Code Generation
+5. Code Generation
 Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation.
-The NLP described in Section IV gives the parameters of the space such as loop order, tiling factor, etc.
+The NLP described in Section 4 gives the parameters of the space such as loop order, tiling factor, etc.
 However, in order to be efficient, the code generation needs some specific rules.
-Communication with Off-Chip Memory  To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler.
+5.1. Communication with Off-Chip Memory
+To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler.
 Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading.
@@
 For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
-Communication between the Fused Task  The communication within the same fused task is not required as they use shared buffers.
+5.2. Communication between the Fused Task
+The communication within the same fused task is not required as they use shared buffers.
 Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
-Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory.
-Data for these tasks resides on-chip.
-If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function.
-The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined.
-Padding is managed at the intra-tile level.
-For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements.
-In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
-Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering.
+5.3. Intra-Task
+Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory.
+Data for these tasks resides entirely on-chip.
+If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined.
+While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
+Padding is handled at the intra-tile level.
+Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage.
+For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
+5.4. Inter-Tile Loop
+For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering.
 Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity.
-Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required.
+Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required.
 This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
-Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
-SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
-VI Evaluation
-VI-A Setup and Experimental Evaluation
-We evaluated our method using kernels from Polybench/C 4.2.1 [56] with medium-sized datasets and single-precision floating-point computations.
+5.5. Concurrent Execution
+Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites. Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
+5.6. SLR Management
+The NLP determines the SLR ID for each task, specifying where it will be executed.
+Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
+5.7. Design Regeneration
+In cases where bitstream generation fails for a given design, we support automatic regeneration of the design.
+Several strategies can be applied, such as tightening resource constraints or reducing the maximum unrolling factor.
+By reducing the available resources, the design becomes smaller, which can help alleviate congestion.
+Thanks to the flexibility of our NLP-based approach, we can retain parts of the previous solution—such as the SLR assignment—and selectively restrict resources only for the specific task or group of tasks responsible for the congestion.
+This is the method we employ to regenerate designs when bitstream generation fails.
+6. Evaluation
+6.1. Setup
+We evaluated our method using kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2024) with medium-sized datasets and single-precision floating-point computations.
 The selected kernels represent both memory-intensive and computation-intensive scenarios.
 Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations.
+Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available.
 NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints.
@@
 RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma.
-The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow [4].
+The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow (AMD/Xilinx, 2024).
 This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory.
 All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
+6.2. Experimental Evaluation
 The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels.
-To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE.
-We execute the designs for these frameworks using identical parameters as outlined in Section VI-A.
-For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory.
-Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers.
-Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications.
+To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, AutoDSE and Stream-HLS.
+For ScaleHLS (Ye et al., 2022), Allo (Chen et al., 2024), and Stream-HLS (Basalama and Cong, 2025), their kernels assume that data is already present in on-chip memory.
+To ensure a fair comparison, we modified their code to incorporate off-chip to on-chip data transfers.
+However, Allo’s 2mm and 3mm kernels already include this transfer mechanism, requiring no modifications.
 We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size.
 We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task.
-For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1].
-Sisyphus generated designs are generated using parameters consistent with the paper [60], and ScaleHLS designs are compiled using their open-source compiler [2].
+For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE (all, 2024).
+Sisyphus generated designs are generated using parameters consistent with the paper (Pouget et al., 2025b).
+We observe some differences in Sisyphus results, as our evaluation is conducted on a different board.
+The design variations between boards may lead to differences in performance.
 For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations.
@@
 If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
-VI-B Comparison
-The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
-RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [60], AutoDSE [67], ScaleHLS [79] (S-HLS), and Allo [14].
+It is worth noting that there is a difference in results between RTL simulation and on-board evaluation.
+This difference is expected, as RTL simulation involves fewer constraints, allowing for more aggressive unrolling and full utilization of available resources.
+In contrast, on-board evaluation is more complex, and excessive unrolling can easily lead to congestion.
+Therefore, we compare RTL simulation results with those of other frameworks to demonstrate our performance improvements.
+More importantly, we show that our approach generates code that can actually be implemented on board.
+6.3. Comparison
+Table 5 presents the RTL simulation results for Prometheus, Sisyphus (Pouget et al., 2025b), AutoDSE (Sohrabizadeh et al., 2021), ScaleHLS (Ye et al., 2022), and Allo (Chen et al., 2024).
 The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels.
+Kernel Ours Sisyphus ScaleHLS Allo AutoDSE Stream-HLS
+2mm 308.38 195.09 37.13 46.58 0.41* 150.27
+3mm 368.36 178.97 43.04 60.40* 1.74* 174.75
+Atax 3.56 2.32 1.58 1.96 1.97* 1.71
+Bicg 15.41 2.32 1.70 14.17 0.99* 1.73
+Gemm 419.14 227.09 40.53 37.50 110.81* 203.48
+Gesummv 10.21 2.28 1.78 8.85 1.98* 1.72
+Mvt 14.65 5.54 7.39 8.77 7.80* 13.31
+Symm 212.20 200.30 0.06 16.72 14.68* N/A
+Syr2k 267.31 149.89 0.08 41.15 12.35* N/A
+Syrk 158.25 105.59 0.27 20.57 23.16* N/A
+Trmm 193.47 166.72 0.07 5.10 0.02* N/A
+PI (Avg) 2.39x 927.20x 8.58x 973.14 3.46x
+PI (gmean) 2.03x 48.03x 4.92x 25.82 2.71x
+Table 5. Throughput Comparison (in GF/s) of PolyBench Kernels Across Frameworks Using RTL Simulation
 Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate.
 The RTL simulation for 3mm with Allo was incomplete after two days.
+The N/A values in the table for Scale-HLS correspond to kernels that have at least one loop with a non-constant trip count.
+Stream-HLS does not handle these cases, so we cannot make a comparison.
 Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks.
-While Sisyphus [60], designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm.
+While Sisyphus (Pouget et al., 2025b), designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm.
 This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications.
 The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
-Kernel Ours Sisyphus S-HLS Allo AutoDSE
-2mm 308.38 195.09 37.13 46.58 0.41*
-3mm 368.36 178.97 43.04 60.40* 1.74*
-Atax 3.56 2.32 1.58 1.96 1.97*
-Bicg 15.41 2.32 1.70 14.17 0.99*
-Gemm 419.14 227.09 40.53 37.50 110.81*
-Gesummv 10.21 2.28 1.78 8.85 1.98*
-Mvt 14.65 5.54 7.39 8.77 7.80*
-Symm 212.20 200.30 0.06 16.72 14.68*
-Syr2k 133.77 75.01 0.04 20.59 12.35*
-Syrk 316.25 211.00 0.54 41.11 23.16*
-Trmm 193.47 166.72 0.07 5.10 0.02*
-PI (Avg) 2.39x 927.20x 8.58x 973.14
-PI (gmean) 2.03x 48.03x 4.92x 25.82
-TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
 Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar.
@@
 Prometheus partially unrolls both loops and pipelines the reduction loop.
-On Board Evaluation  Table III presents the results for the on-board evaluation.
+On Board Evaluation  Table 6 presents the results for the on-board evaluation.
 The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs.
@@
 Bicg 0.14 4.34 1226 241 291 380 177
-TABLE III: On-board evaluation comparison
+Table 6. On-Board Evaluation of Performance and Resource Utilization Across Frameworks and SLR Configurations
 Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE.
@@
 When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
-Table IV shows the parameters found by the NLP.
-We use the same name of the iterator as Polybench 4.2.1 [56] code.
+Table 7 shows the parameters found by the NLP.
+We use the same name of the iterator as Polybench 4.2.1 (Pouchet and Yuki, 2024) code.
 $Si$ represent the statement in position $i$ in the code.
-The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
+The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
 Permutations are evident in the implementations of 3mm, Atax, and Bicg.
@@
 Bicg FT0: S1, S2, FT1: S0, S3 FT0: j,i FT1: i,j s: $16$, A (FT0): $416\times 16$, r: $416$, q: $10$, A (FT1): $10\times 400$, p: $400$,
-TABLE IV: Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table III for 1 SLR
+Table 7. Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table 6 for 1 SLR
 As other frameworks cannot generate bitstreams for 3 SLRs without human intervention, we compare results for 1 and 3 SLRs using our framework.
@@
 However, for atax and bicg, where the bottleneck is memory transfer between off-chip and on-chip rather than parallelism, the improvement is negligible.
-VII Related Work
-Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [22], synchronous dataflow languages [42, 11], and for FPGA applications [53, 12, 3, 25, 10, 5, 75, 14].
-DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention.
-Flower [5] automates FPGA dataflow development but limits parallelism.
-Frameworks like [14, 75], built on HeteroCL [38], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization.
-Systolic arrays [71, 20, 33, 39, 9] offer efficient computation for specific patterns but lack generalization.
-Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
-Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [58, 59, 60], bottleneck DSE [67], and GNN-based latency and resource estimation [66, 64, 65, 8, 68, 73, 74].
+6.4. Scalability
+Our NLP solver includes the option to set a timeout, returning the best design found so far without guaranteeing optimality.
+This feature enables efficient exploration of the solution space while ensuring adherence to time constraints when necessary.
+When analyzing the time required to solve the Non-Linear Programming (NLP) problem in Sisyphus, we observe that our framework, Prometheus, achieves similar solution times for most benchmarks.
+However, there is a notable exception: the 3mm benchmark, which times out after 4 hours in Sisyphus, while Prometheus successfully finds a solution.
+This efficiency stems from Prometheus’ ability to explore a larger optimization space while still maintaining fast solution times.
+The key reason is that all optimization techniques are seamlessly integrated within the design space.
+This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity.
+As a result, even with a broader search space, Prometheus efficiently converges to a solution in 21.37 seconds.
+A detailed comparison of solution times can be found in Table 8, which highlights the impact of our approach in managing the design space effectively.
+Benchmark Sisyphus Prometheus
+2mm 22.37 15.98
+3mm 14400.08 21.37
+Atax 3.18 7.03
+Bicg 3.17 5.09
+Gemm 2.19 4.31
+Gesummv 1.57 8.07
+Mvt 1.71 1.26
+Symm 7.13 11.46
+Syr2k 5.29 6.07
+Syrk 2.56 3.94
+Trmm 4.34 5.96
+Average 1313.96 8.23
+Geo Mean 7.95 6.50
+Table 8. Time Required (in seconds) for the NLP Solver to Find an Optimal Solution: Sisyphus vs. Prometheus Across Benchmarks (Timeout Set at 14,400s)
+7. Related Work
+Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024).
+DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention.
+Stream-HLS (Basalama and Cong, 2025) automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory.
+Additionally, it offers very limited opportunities for parallelism.
+Flower (Amiri et al., 2021) automates FPGA dataflow development but limits parallelism.
+Frameworks like (Chen et al., 2024; Xiang et al., 2022), built on HeteroCL (Lai et al., 2019), optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization.
+Systolic arrays (Wang et al., 2021; Cong and Wang, 2018; Jia et al., 2020; Lai et al., 2020; Basalama et al., 2023) offer efficient computation for specific patterns but lack generalization.
+Application-specific frameworks (Chi et al., 2018; Hegarty et al., 2014; Natale et al., 2016; de Fine Licht et al., 2021; Singh et al., 2020; Denzler et al., 2023; Korol et al., 2022; Bacis et al., 2017) demonstrate dataflow advantages but do not generalize across domains.
+RapidStream-TAPA (Guo et al., 2023a, b) enhances the performance of dataflow designs and automates SLR placement.
+However, it requires an optimized kernel with a dataflow structure as input.
+Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b).
 However, these approaches lack effective integration of dataflow optimization techniques.
-Code Transformation  Code transformation has been explored for CPUs [13, 55, 37, 7], GPUs [70], and FPGAs [57, 84, 83, 44, 48, 47, 46, 17].
-Pluto [13] minimizes communication and improves locality but can limit parallelism.
-FPGA-specific adaptations [57] leverage FIFOs for overlapping communication and computation but are restricted in parallelism.
-Recent works [84, 83] selectively use Pluto for latency minimization while avoiding non-HLS-friendly code.
-While [44, 48, 47, 46, 17] focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives.
-The [79, 38, 31, 14, 78] compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties.
-The paper [60] described in Section II generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution.
+Code Transformation  Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018).
+Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism.
+FPGA-specific adaptations (Pouchet et al., 2013b) leverage FIFOs for overlapping communication and computation but are restricted in parallelism.
+Recent works (Zhao et al., 2022; Zhao and Cheng, 2021) selectively use Pluto for latency minimization while avoiding non-HLS-friendly code.
+While (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives.
+The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties.
+The paper (Pouget et al., 2025b) described in Section 2 generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution.
 Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
 Tiling and Padding  Tiling is essential for balancing computation and communication.
-While prior works [49, 45] use cost models to minimize communication, our approach extends this to reduce overall latency.
-Techniques like NLP-based tiling [45] focus on CPUs, while Wedler [61] optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts.
-Padding is well-studied for reducing cache misses [30, 52] and improving memory transfers [72], but its use for varying unroll factors on FPGAs remains underexplored.
-VIII Conclusion
-In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework.
-By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints.
-Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.
+While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency.
+Techniques like NLP-based tiling (Li et al., 2021) focus on CPUs, while Wedler (Shi et al., 2023) optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts.
+Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
+8. Conclusion
+In this work, we introduced Prometheus, a holistic optimization framework that unifies loop transformations, task concurrency, computation-communication overlap, and hardware-aware scheduling for FPGA accelerators.
+By formulating the optimization process as a Non-Linear Programming (NLP) problem, Prometheus enables a structured exploration of the vast design space while considering hardware constraints, memory bandwidth, and task parallelism.
+Our framework addresses key limitations in existing methodologies, which often optimize only isolated aspects of FPGA acceleration.
+Prometheus surpasses these approaches by integrating SLR-aware scheduling, dynamic memory management, and hybrid execution models that effectively balance dataflow streaming and shared buffering.
+This enables more efficient utilization of FPGA resources, leading to significant improvements in latency, throughput, and scalability.
+Through extensive performance evaluations on computation-bound kernels, we demonstrated that Prometheus outperforms state-of-the-art frameworks.
+Furthermore, Prometheus’ automatic design space exploration significantly reduces the manual effort required in FPGA development.
+By automatically generating HLS-C++ code, OpenCL host code, and FPGA bitstreams, the framework streamlines the deployment process, making FPGA acceleration more accessible to a broader range of applications.
+In summary, Prometheus provides an innovative and effective approach to FPGA optimization, delivering high-performance solutions while minimizing the complexity of hardware design.
+Acknowledgments
+This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.

### v5

@@
 The results showcase its superior performance compared to state-of-the-art FPGA optimization frameworks, highlighting its effectiveness in delivering high QoR while reducing manual tuning efforts.
-High-Level Synthesis, NonLinear Programming, Compiler
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
+High-Level Synthesis, Non-Linear Programming, Compiler
+copyright: ccjournal: TODAESdoi: 10.1145/3769307ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
 1. Introduction
@@
 However, achieving a good Quality of Result (QoR) on FPGAs remains a complex and resource-intensive process, requiring careful tuning of computation, memory access, and parallelism strategies.
-To address this complexity, High-Level Synthesis (HLS) tools and source-to-source compilers (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025) have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python.
+To address this complexity, High-Level Synthesis (HLS) tools (Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023) and source-to-source compilers e.g,  (Ye et al., 2022; Lai et al., 2019; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025), have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python.
 These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution.
@@
 A more adaptive approach is needed to orchestrate data movement in parallel with execution, reducing overall memory stalls.
-Scalability and Multi-SLR Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability.
-However, most frameworks, including Merlin-based tools (e.g., AutoDSE, HARP, NLP-DSE) and Sisyphus, restrict designs to a single SLR, leading to underutilization of FPGA resources.
+Scalability and Multi-Super Logic Regions Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability.
+However, most frameworks, including Merlin-based tools restrict designs to a single SLR, leading to underutilization of FPGA resources.
 While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity.
@@
 1.2. Contributions and Prometheus Method
-This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology.
-Unlike existing tools that optimize isolated aspects of FPGA acceleration, Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling.
+This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology specialized for affine programs (Lattner et al., 2021).
+Specifically, we operate on affine loop nests where statements can be distributed in multiple loops (i.e.
+loops are permutable (Pouchet et al., 2011)).
+Such programs encompass typical computational pattern in dense linear algebra, data mining, image processing, etc.
+as exemplified in the PolyBench benchmarking suite (Polybench, [n.
+d.]; Abella-González et al., 2021).
+Unlike existing tools that optimize isolated aspects of FPGA acceleration such as AutoDSE (Sohrabizadeh et al., 2021), HARP (Sohrabizadeh et al., 2023), NLP-DSE (Pouget et al., 2025a), Sisyphus (Pouget et al., 2025b) and Stream-HLS (Basalama and Cong, 2025), Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling.
 It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
@@
 Once the best configuration is identified, Prometheus automatically generates the optimized design.
-Finally, the system produces an OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
+Finally, the system produces OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
 The key contributions of this work are:
-  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurency and computation-communication overlap while considering interdependencies between computation and data movement.
+  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurrency and computation-communication overlap while considering interdependencies between computation and data movement.
   • Hybrid execution model that dynamically selects between shared buffering and dataflow streaming to maximize parallelism and efficient memory access.
-  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies for globally theoretical optimal performance.
+  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies to achieve a globally optimal theoretical performance.
   • SLR-aware scheduling and multi-SLR partitioning to balance routing complexity and FPGA resource utilization, overcoming single-SLR limitations in prior works.
-  • End-to-end compilation and automation that generates optimized HLS-C++ code, OpenCL host code, and FPGA bitstreams with minimal manual intervention.
+  • End-to-end compilation and automation, generating optimized HLS-C++ code—that is, standard C++ source code annotated with AMD/Vitis-specific high-level synthesis (HLS) pragmas to guide hardware generation—alongside OpenCL host code and FPGA bitstreams, all with minimal manual intervention.
   • Comprehensive performance evaluation, demonstrating superior QoR compared to AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), ScaleHLS (Ye et al., 2022), Stream-HLS (Basalama and Cong, 2025) and Allo (Chen et al., 2024).
+Our framework, Prometheus, is open source and available at https://github.com/UCLA-VAST/Prometheus.
 2. Background and Motivation
@@
 AutoDSE (Sohrabizadeh et al., 2021) HARP (Sohrabizadeh et al., 2023) PolyOpt-HLS (Pouchet et al., 2013a) ScaleHLS (Ye et al., 2022) / POM (Zhang et al., 2024) HeteroCL (Lai et al., 2019) / Allo (Chen et al., 2024) Stream HLS (Basalama and Cong, 2025) NLP-DSE (Pouget et al., 2025a) Sisyphus (Pouget et al., 2025b) Prometheus
-Pragma Insertion Limit.
+Pragma Insertion Limit
 Code Transformation (Tiling) Limit Limit
@@
 Code transformation + pragma insertion (unified)
-Task Concurrency Limit.
+Task Concurrency Limit
 Dataflow
-Computation-Communication Overlap Limit.
+Computation-Communication Overlap Limit
 Data Packing
@@
 Padding (Computation)
-Hardware Aware Limit.
-Hardware-feasible Limit. Limit. Limit. Limit.
+Hardware Aware Limit
+Hardware-feasible Limit Limit Limit Limit
 Management of Off-Chip and On-Chip Memory Transfers
@@
 By enforcing a producer-consumer model, dataflow scheduling enables each task to process data as soon as it becomes available, rather than waiting for an entire stage to complete, which is particularly beneficial.
-For computational kernels such as 3mm (Listing 3), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced.
+For computational kernels such as 3mm (Listing 4), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced.
 This overlapping of execution helps maximize throughput and minimize idle time for computing units.
@@
 To further enhance intra-task parallelism, alternative approaches must be explored.
-Stream-HLS (Basalama and Cong, 2025) attempts to address this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip.
+Stream-HLS (Basalama and Cong, 2025) addressed this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip.
 While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory.
@@
 By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance.
-It is particularly beneficial in applications which are computation bound.
+It is particularly beneficial in applications that are computation-bound.
 However, shared buffering presents challenges in maintaining concurrency.
@@
 Data packing and padding optimize memory alignment, reducing the number of required memory cycles.
-For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles, making efficient packing crucial for minimizing overhead.
+For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles – compared to 216 cycles without packing.
+This highlights the importance of efficient data packing to minimize overhead.
+However, to enable such packing, the transfer vector size must evenly divide the total array size.
 To fully exploit data packing, padding must be considered to enable even faster data transfers.
@@
 However, padding is not a free optimization, as increasing it also increases the amount of data that needs to be transferred.
-For example, in the code presented in Listings 1 and 2, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth.
+Figure 1. Illustration of Padding Strategy to Align Array Size for Efficient Data Packing and Memory Transfers on AMD FPGAs
+1// Original loop without padding
+2for (int i = 0; i < 190; i++){
+3#pragma HLS unroll factor=UF
+4// Possible UF values: 1, 2, 5, 10, 19, 38, 95, 190
+5 A[i] += 1;
+6}
+7
+8// Padded loop
+9for (int i = 0; i < 190 + 2; i++){ // Padding of 2
+10#pragma HLS unroll factor=UF
+11// Possible UF values: 1, 2, 3, 4, 6, 8, 12, 16, 24, 32, 48, 64, 96, 192
+12 A[i] += 1;
+13}
+Listing 1: Effect of Padding on the Space of Unroll Factors for Computation
+Figure 1 illustrates the selection of the padding value $P$ for an array $A$ of size $N$.
+By leveraging data packing (up to 512 bits), multiple elements can be transferred per cycle.
+However, due to constraints in AMD compilers and FPGA architectures, the total transfer size must be divisible by a power of two.
+Padding is therefore introduced to align the data size accordingly and enable more efficient transfers.
+In the code presented in Listings 2 and 3, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth.
 Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$.
@@
 Padding for Computation
-Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 1 to Listing 2.
-We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra ressource to only compute a partial tile which would correspond to the execution of the last tile which is not complete.
+Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 2 to Listing 3.
+We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra resource to only compute a partial tile which would correspond to the execution of the last tile which is not complete.
 For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
-But we restrict the unroll factor to be divisible of the trip count we restrict considerably the space of possibilities.
-So the solution is to pad which allow to considerably increase the space of possibilities.
-Hence with a trip count of $190$ the possible unroll factor are include in $UF\in\{1,2,5,10,19,38,95,190\}$ but with padding of 2 (a trip count of 192) the space become $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}$.
-And of course all pading can be considered.
-Padding enables finer control over unrolling and resource utilization, allowing for more flexible and efficient hardware implementation while optimizing computational parallelism.
+If the unroll factor is restricted to be a divisor of the loop trip count, the design space becomes significantly limited.
+Padding can be used to adjust the trip count, thereby expanding the set of valid unroll factors to better match available hardware resources.
+As shown in Listing 1, a loop with a trip count of 190 permits only a limited set of unroll factors: $UF\in\{1,2,5,10,19,38,95,190\}.$ However, by padding the loop to a trip count of 192, the space of legal unroll factors becomes: $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}.$ This increased flexibility allows finer control over unrolling and resource utilization, enabling more efficient hardware implementations and better exploitation of computational parallelism.
 1for (i = 0; i < I; i++)
@@
 4 C[i][j] += A[i][k] * B[k][j];
-Listing 1: Baseline Implementation of Matrix Multiplication in C.
+Listing 2: Baseline Implementation of Matrix Multiplication in C.
 This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
@@
 10 C[i0*I1+i1][j0*J1+j1] += A[i0*I1+i1][k0*K1+k1]*B[k0*K1+k1][j0*J1+j1];
-Listing 2: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation.
+Listing 3: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation.
 This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
@@
 Refer to Table 1 for a detailed comparison of these approaches.
-A common oversimplification in many frameworks is the assumption that all data resides on-chip.
+A common oversimplification in many frameworks is the assumption that all data resides on-chip e.g., (Basalama and Cong, 2025; Ye et al., 2022).
 While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware.
@@
 It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication.
-As illustrated in Listing 3, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap.
+The 3mm kernel computes: G = (A × B) × (C × D) where intermediate matrices E and F store the results of the two first multiplications.
+As illustrated in Listing 4, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap.
 However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
@@
 While they employ a cost model to minimize computation latency, they still rely on exhaustive enumeration of possible configurations.
-Allo (Chen et al., 2024) aims to reduce development time and enhance design quality.
+Allo (Chen et al., 2024) aims to reduce development time and enhance design quality through a composable programming model for hardware accelerator design.
+By decoupling algorithm specification from hardware customization and enabling modular schedule composition, Allo supports holistic dataflow optimization and verifiable transformations for large-scale, multi-kernel designs.
 However, it still requires significant manual intervention from an expert, limiting its usability for non-specialists.
@@
 Even if the bitstream is successfully generated, timing violations often degrade performance.
-1for (i = 0; i < 180; i++) // MM 1
+1for (i = 0; i < 180; i++) // $M$ 1
 2 for (j = 0; j < 190; j++) {
@@
 5 E[i][j] += A[i][k]*B[k][j];}// S1
-6for (i = 0; i < 190; i++) // MM 2
+6for (i = 0; i < 190; i++) // $M$ 2
 7 for (j = 0; j < 210; j++) {
@@
 10 F[i][j] += C[i][k]*D[k][j];}// S3
-11for (i = 0; i < 180; i++) // MM 3
+11for (i = 0; i < 180; i++) // $M$ 3
 12 for (j = 0; j < 210; j++) {
@@
 15 G[i][j] += E[i][k]*F[k][j];}// S5
-Listing 3: Reference Implementation of the 3mm Kernel from PolyBench.
-This listing shows the baseline version of the 3mm benchmark, comprising three matrix multiplication operations.
-The first two multiplications are independent and produce intermediate results that serve as inputs to the final matrix multiplication.
+Listing 4: Reference Implementation of the 3mm Kernel from PolyBench.
+The 3mm kernel computes: G = (A × B) × (C × D) where intermediate matrices E and F store the results of the two first multiplications.
+1for (i = 0; i < 180; i++) // Task 0
+2 for (j = 0; j < 190; j++)
+3 E[i][j] = 0.0; // S0
+4for (i = 0; i < 180; i++) // Task 1
+5 for (j = 0; j < 190; j++)
+6 for (k = 0; k < 200; ++k)
+7 E[i][j] += A[i][k]*B[k][j];}// S1
+8for (i = 0; i < 190; i++) // Task 2
+9 for (j = 0; j < 210; j++)
+10 F[i][j] = 0.0; // S2
+11for (i = 0; i < 190; i++) // Task 3
+12 for (j = 0; j < 210; j++)
+13 for (k = 0; k < 220; ++k)
+14 F[i][j] += C[i][k]*D[k][j];}// S3
+15for (i = 0; i < 180; i++) // Task 4
+16 for (j = 0; j < 210; j++)
+17 G[i][j] = 0.0; // S4
+18for (i = 0; i < 180; i++) // Task 5
+19 for (j = 0; j < 210; j++)
+20 for (k = 0; k < 190; ++k)
+21 G[i][j] += E[i][k]*F[k][j];}// S5
+Listing 5: Direct Transformation of Listing 4 Where Each Loop Body Becomes a Separate Task
 2.4. Overview of Prometheus
@@
 DSP Budget Maximum DSP slices per SLR; used to limit total DSP usage by pipelined/unrolled tasks.
-On-Chip Memory Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
+On-Chip Memory Available Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
 Table 2.
@@
 The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information.
-This extraction is performed using PoCC (Pouchet, 2023), which provides all necessary details for analysis.
+This extraction is performed using PoCC (PoCC, [n.
+d.]), which provides all necessary details for analysis.
 Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem.
@@
 These variables and constraints jointly define the feasible design space that the NLP-based optimization engine systematically explores to generate high-performance, hardware-aware FPGA implementations.
-The pseudo-code in Listing 4 demonstrates how Prometheus processes the 3mm kernel (Listing 3).
+The pseudo-code in Listing 6 demonstrates how Prometheus processes the 3mm kernel (Listing 4).
 Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
-Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 3. An example of such a task is illustrated in Listing 5.
+Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 4. An example of such a task is illustrated in Listing 7.
 Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
-Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 3), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
+Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 4), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding.
 The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
@@
 35 store_G(G);
-Listing 4: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus.
+Listing 6: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus.
 This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.
@@
 If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
+Figure 3 shows the task graph of the 3mm kernel (Listing 5), where each loop body corresponds to a distinct computation task.
+T0T1T2T3T4T5
+Figure 3. Dataflow graph of the 3mm kernel. Nodes represent computation tasks; edges denote data dependencies.
 Prometheus’s solution space includes:
@@
 Tile sizes are consistent within a fused task but vary between tasks.
-In Listing 4, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
+In Listing 6, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
 3.3. Fine-grained Parallelism
@@
 For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction.
-For instance, in Listing 4, the inter-tile loops on lines 3 and 5 iterate over task0 and task1.
+For instance, in Listing 6, the inter-tile loops on lines 3 and 5 iterate over task0 and task1.
 The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip.
@@
 Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency.
-For instance, in Listing 4, additions take 3 cycles, resulting in $II=3$.
+For instance, in Listing 6, additions take 3 cycles, resulting in $II=3$.
 3.4. Loop Order
@@
 Section 4).
-As shown in Listing 4, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 3.
+As shown in Listing 6, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 4.
 3.5. Automatic Overlapping of Communication and Computation
@@
 Constants Description
-$II_{l}$ II of the loop $l$
-$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
+$I_{l} II$ of the loop $l$
+$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without (parpar) and with (redred) dependencies of the statement $s$
 $TC^{l}_{ori}$ Original Trip Count of the loop $l$
@@
 $N_{FT}$ Number of fused task
-$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
-$DSP$ Number of DSP available for the FPGA used
+$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation opop
+DSPDSP Number of DSP available for the FPGA used
 $max_{part}$ Maximum array partitioning
-$SLR$ Number of SLR available for the FPGA used
+SLRSLR Number of SLR available for the FPGA used
 Variables Description
@@
 4.1. Constraints
-We now describe the constraints by using the code of Listings 3, 4 and 5.
+We now describe the constraints by using the code of Listings 4, 6 and 7.
 1for (int j1 = 0; j1 < 32; j1++)
@@
 8 F[i1][j1] += C[i][k] * D[k][j1];}
-Listing 5: Implementation of Task 3 in the 3mm Kernel (from Listing 4).
+Listing 7: Implementation of Task 3 in the 3mm Kernel (from Listing 6).
 This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism.
@@
 The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver.
-For instance, in Listing 5, for the array $F$, the loop $j$ (line 7 in Listing 3) has been split into $j0$ (line 14 in Listing 4) and $j1$ (intra-tile).
+For instance, in Listing 7, for the array $F$, the loop $j$ (line 7 in Listing 4) has been split into $j0$ (line 14 in Listing 6) and $j1$ (intra-tile).
 The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
@@
 Equation 6 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
-The array $E$, defined on line 24 in Listing 4, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops).
+The array $E$, defined on line 24 in Listing 6, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops).
 However, it is transferred under the loop $i0$, so $t_{E,i0}=1$.
@@
 We cannot transfer E under loop $i0$ if E is defined under $k0$.
-Similarly, in Listing 4, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
+Similarly, in Listing 6, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
 (5) $\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$
@@
 Equation 9 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
-Array D in Listing 5 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$).
+Array D in Listing 7 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$).
 Therefore, the total number of partitions needed is $3\times 32=96$.
@@
 4.2. Objective Function
-The objective function models the dependency graph of the fused task.
-Independent tasks execute simultaneously, with latency set by the longest task.
-Dependent tasks overlap partially, with the second task starting once it receives enough data from the first.
-The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts.
-For our 3mm example, the graph is shown in Figure 2, and the corresponding latency is:
-$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
-PrometheusT0T1T2T3T4T5
-Figure 2. Maximally distributed dataflow graph of the 3mm kernel, where each node represents a task and each edge indicates a data dependency between tasks.
-The latency of the intra-task for a statement $s$ is determine similarly to (Pouget et al., 2025a, 2024, b) with Equation 12.
-And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
-(12) $Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$
-(13) $Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$
-For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication.
-The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
-(14) $Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$
-And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
+The objective of our framework is to minimize the total latency of the design, which is modeled as a directed acyclic graph (DAG).
+In this dataflow graph, each node represents a computation task (or fused task), and edges represent data dependencies between tasks.
+To compute the overall latency, we define the latency of each task $T\in\mathcal{T}$ as the sum of:
+  • the time at which it becomes ready to start (determined by its dependencies)
+  • and the duration of the task itself
+We denote:
+  • $\text{Lat}(T)$: the global latency contribution of task $T$, i.e., the time at which $T$ finishes,
+  • $\text{Lat}_{\text{task}}(T)$: the execution time (duration) of task $T$,
+  • $\text{pred}(T)$: the set of immediate predecessor tasks of $T$,
+  • $\text{shift}_{T_{i},T}$: the number of cycles after which $T$ can start once $T_{i}$ has begun (e.g., due to pipelined data production).
+Then the latency of each task is recursively defined as:
+(12) $\text{Lat}(T)=\max_{T_{i}\in\text{pred}(T)}\left[\text{Lat}(T_{i})+\text{shift}_{T_{i},T}\right]+\text{Lat}_{\text{task}}(T)$
+The overall latency of the design is defined as the latest finish time among all sink tasks (i.e., tasks without successors):
+(13) $\text{Lat}_{\text{total}}=\max_{T\in\mathcal{S}}\left[\text{Lat}(T)\right]$
+where $\mathcal{S}\subseteq\mathcal{T}$ is the set of sink tasks in the graph.
+For the graph in Figure 3, the latency can be expressed step-by-step as:
+$\left.\begin{aligned} &\text{Lat}(T_{1})=\text{Lat}(T_{0})+\text{shift}_{T_{0},T_{1}}+\text{Lat}_{\text{task}}(T_{1})\\ &\text{Lat}(T_{3})=\text{Lat}(T_{2})+\text{shift}_{T_{2},T_{3}}+\text{Lat}_{\text{task}}(T_{3})\\ &\text{Lat}(T_{5})=\max\left(\text{Lat}(T_{1})+\text{shift}_{T_{1},T_{5}},\;\text{Lat}(T_{3})+\text{shift}_{T_{3},T_{5}},\;\text{Lat}(T_{4})+\text{shift}_{T_{4},T_{5}}\right)+\text{Lat}_{\text{task}}(T_{5})\end{aligned}\right$.
+In the next section, we explain how to compute $\text{Lat}_{\text{task}}(T)$ for each task using intra-task and inter-tile modeling.
+4.2.1. Intra-Task and Multi-Level Latency Modeling
+Each task $T$ is composed of computation over a tile of data and may be enclosed in multiple inter-tile loop levels (e.g., tiling for cache reuse).
+At each level, we must account for both computation and communication (i.e., data transfers).
+We model the execution as a combination of:
+  • Load: reading data from off-chip memory (or previous task) into on-chip buffers,
+  • Compute: performing the actual tile computation,
+  • Store: writing the result back to off-chip memory (or to a subsequent task).
+To enable communication/computation overlap, we use double- or triple-buffering depending on the number of streams.
+At each inter-tile level $n$, the total latency is computed as the maximum of the three components, accounting for pipeline shifts and buffer reuse.
+Level-Based Latency Recursion.
+Let:
+  • $\text{Lat}_{n+1}$: latency of executing the inner level $n+1$,
+  • $f_{a,n}$: number of bytes transferred for array $a$ at level $n$,
+  • $BS_{a}$: memory bandwidth for array $a$,
+  • $\alpha$: overlap factor (1 = load or store only, 2 = both).
+Then, the latency at level $n$ is:
+(14) $\text{Lat}_{n}=\max\left(\text{Lat}_{n+1},\frac{f_{a,n}}{BS_{a}}\right)+\text{Lat}_{n+1}+\alpha\cdot\frac{f_{a,n}}{BS_{a}}$
+This formula captures both the initialization overhead and the overlap between communication and computation.
+Base Case: Intra-Task Latency.
+The base latency $\text{Lat}_{n+1}$ is the latency of the innermost computation tile, computed as in (Pouget et al., 2025a, b).
+(15) $\text{Lat}_{\text{intra}}=IL_{\text{par}}+IL_{\text{seq}}\cdot\log_{2}\left(\prod_{l\in\mathcal{L}^{\text{red}}_{s}}TC^{l}_{\text{intra}}\right)$
+If the computation is invoked across multiple tiles in a pipelined loop, the full latency becomes:
+(16) $\text{Lat}_{\text{task}}=\text{Lat}_{\text{intra}}+II\cdot\left(\prod_{l\in\mathcal{L}^{\text{red}}_{s}}TC^{l}_{\text{inter}}-1\right)$
+Final Task Latency.
+The full latency of a task $T$, denoted $\text{Lat}_{\text{task}}(T)$, is the latency at the outermost loop level $n=0$, computed recursively using Equation 14, where the base is given by Equation 16.
 5. Code Generation
@@
 For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
+Listing 8 shows the load and read functions, which first transfer data from off-chip memory to an on-chip FIFO, and then read from the FIFO into the local buffer allocated for Task 1.
+1void load_vA_for_task1(hls::stream<float16> &fifo_A_from_off_chip_to_S1,
+2 float16 vA[2340]) {
+3#pragma HLS inline off
+4 for (int i = 0; i < 2340; i++) {
+5#pragma HLS pipeline II = 1
+6 fifo_A_from_off_chip_to_S1.write(vA[i]);
+7 }
+8}
+9
+10void read_A_FT0(float A[10][204],
+11 hls::stream<float16> &fifo_A_from_off_chip_to_S1, int i0) {
+12#pragma HLS inline off
+13 if (i0 >= 18) {
+14 return;
+15 }
+16 for (int d0 = 0; d0 < 10; d0++) {
+17 for (int d1 = 0; d1 < 204; d1 += 16) {
+18#pragma HLS pipeline II = 1
+19 float16 tmp_fifo = fifo_A_from_off_chip_to_S1.read();
+20 for (int j = 0; j < 16; j++){
+21#pragma HLS unroll
+22 if (d1 + j < 204)
+23 A[d0][d1 + j] = tmp_fifo[j];
+24 }
+25 }
+26 }
+27}
+Listing 8: Memory transfer for array A in the 3mm example (Listing 4). The data is first loaded from off-chip memory into a FIFO, then read from the FIFO to populate the on-chip buffer
 5.2. Communication between the Fused Task
@@
 Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
+Listing 9 illustrates an example where the write function transfers data produced by Tasks 0 and 1 into a FIFO, which is then consumed by Task 5.
+As shown, the on-chip buffers used by Tasks 0 and 1 differ from those used by Task 5, allowing greater flexibility in choosing tile sizes for each task independently.
+1void write_E_FT0(float E[10][32],
+2 hls::stream<float16> &fifo_E_from_task1_to_task5, int j0,
+3 int i0) {
+4#pragma HLS inline off
+5 if (j0 < 0 $|$ i0 < 0) {
+6 return;
+7 }
+8 for (int d0 = 0; d0 < 10; d0++) {
+9 for (int d1 = 0; d1 < 32; d1 += 16) {
+10#pragma HLS pipeline II = 1
+11 float16 tmp_fifo;
+12 for (j = 0; j < 16; j++){
+13#pragma HLS unroll
+14 tmp_fifo[j] = E[d0][d1 + j];
+15 }
+16 fifo_E_from_task1_to_task5.write(tmp_fifo);
+17 }
+18 }
+19}
+20
+21void read_E_FT2(float E[180][192],
+22 hls::stream<float16> &fifo_E_from_task1_to_task5, int i0,
+23 int j0) {
+24#pragma HLS inline off
+25 if (j0 > 0 $|$ i0 >= 18) {
+26 return;
+27 }
+28 for (int d1_0 = 0; d1_0 < 6; d1_0++) {
+29 for (int d0 = 0; d0 < 10; d0++) {
+30 for (int d1_1 = 0; d1_1 < 32; d1_1 += 16) {
+31 int d1 = d1_0 * 32 + d1_1;
+32 float16 tmp_fifo = fifo_E_from_task1_to_task5.read();
+33 for (j = 0; j < 16; j++){
+34#pragma HLS unroll
+35 if (d1 + j < 192)
+36 E[d0 + i0 * 10][d1 + 0 + j] = tmp_fifo[j];
+37 }
+38 }
+39 }
+40 }
+41}
+Listing 9: FIFO-Based Inter-Task Communication for Intermediate Matrix E: The first function writes data from the local buffer E into a FIFO, while the second reads from the FIFO to populate the next computation stage’s on-chip buffer.
 5.3. Intra-Task
@@
 Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available.
+We include a comparison with Sisyphus using the n-madd kernels, where $n$ denotes the number of matrix additions.
+In the case of 2-madd, the result of the first addition is used as the input to the second.
+For 3-madd, the results of the first two additions are both used as inputs to the final addition.
+Table 5 lists the benchmark kernels used in our evaluation.
+For each kernel, we report the computational complexity (in terms of the number of operations) and the memory complexity (in terms of footprint of the input/output data).
+Additionally, we include the data reuse order, which reflects the approximate number of times one data element is reused across different computations within the kernel.
+Kernels with reuse on the order of $\mathcal{O}({N})$, where $N$ denotes the problem size (e.g., the number of rows or columns in an $N\times N$ input matrix), are typically considered compute-bound as their theoretical arithmetic intensity (assuming perfect reuse on-chip) is $\mathcal{O}({N})$.
+In contrast, kernels with reuse on the order of $\mathcal{O}({1})$ are generally classified as memory-bound.
+In practice, compute-bound kernels require very careful bufferization on-chip to minimize off-chip communications and achieve the theoretical arithmetic intensity.
+The final column (Communication Between Tasks) represents the number of data elements that are transferred between tasks in the dataflow design, excluding any initial input loading or setup overhead.
+Benchmark Description Ops      Complexity Mem      Complexity Reuse Comm. Between Tasks
+bicg BiCG sub-kernel of BiCGStab solver $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
+madd Matrix add. (C = A + B) $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
+mvt Matrix Vector product and Transpose $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ 0
+atax Matrix transpose and vector mult. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ NN
+gesummv Scalar, vector and matrix mult. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $2N$
+2-madd 2 Matrix add. (D = (A + B) + C) $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $N^{2}$
+[…]
+gemver Vector mult. and matrix add. $\mathcal{O}(N^{2})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(1)$ $2N^{2}+2N$
+2mm 2 Matrix Mult. ($\alpha ABC+\beta D$) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
+gemm Matrix-multiply ($C=\alpha AB+\beta C$) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
+syr2k Symmetric rank-2k update $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
+syrk Symmetric rank-k update $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
+trmm Triangular matrix-mult. $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $N^{2}$
+[…]
+symm Symmetric matrix-mult. $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ $\mathcal{O}(N)$ $2N^{2}$
+Table 5. Benchmark Kernels Used for Evaluation: Computational Complexity, Memory Requirements, Data Reuse, and Inter-Task Communication Analysis
 NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints.
@@
 AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs.
-If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
+If congestion persists within an SLR, we manually adjust the NLP constraints based on the specific congestion issue and regenerate the HLS-C++ code, a process that typically takes only a few minutes.
+As our NLP model includes many constraints, both global and per SLR.
+If congestion occurs, we can address it by adjusting the relevant constraint.
+For example, if the issue is due to resource overutilization, we can retain the previous NLP solution and simply tighten the resource utilization constraint for the specific SLR.
+It is worth noting that different regeneration strategies are possible, depending on whether we prioritize achieving a high-quality result (QoR) or obtaining a solution quickly.
+For now, this decision is left to the user.
+This step is currently performed manually to better understand the root causes of congestion, but it could be automated through analysis of the compiler-generated log files.
 It is worth noting that there is a difference in results between RTL simulation and on-board evaluation.
@@
 Trmm 193.47 166.72 0.07 5.10 0.02* N/A
-PI (Avg) 2.39x 927.20x 8.58x 973.14 3.46x
-PI (gmean) 2.03x 48.03x 4.92x 25.82 2.71x
+PI (Avg) 1.00x 2.39x 927.20x 8.58x 973.14x 3.46x
+PI (gmean) 1.00x 2.03x 48.03x 4.92x 25.82x 2.71x
 Table 6. Throughput Comparison (in GF/s) of PolyBench Kernels Across Frameworks Using RTL Simulation
@@
 Prometheus partially unrolls both loops and pipelines the reduction loop.
+Sisyphus Prometheus
+Kernel GF/s BRAM (%) DSP (%) FF (%) LUT (%) GF/s BRAM (%) DSP (%) FF (%) LUT (%)
+Madd 1.96 1 29 22 38 2.71 3 7 10 11
+2-madd 3.89 1 58 59 91 8.99 6 14 21 24
+3-madd 3.91 2 58 75 117 14.00 9 21 31 37
+2mm 195.09 3 73 49 72 308.38 19 71 41 63
+3mm 178.97 16 98 86 122 368.36 20 83 52 85
+Gemm 227.09 6 51 42 72 419.14 18 53 33 45
+Gemver 17.93 24 116 56 68 22.51 31 53 34 54
+Mvt 5.54 23 2 6 10 14.65 6 22 16 51
+Table 7. RTL Evaluation of Performance and Resource Utilization for Sisyphus and Prometheus
+Table 7 presents a comparison between Prometheus and Sisyphus on additional kernels.
+As shown, Prometheus generally achieves higher throughput and better overall resource utilization, except for BRAM usage, which is higher due to the use of double buffering.
+Notably, the 3-madd kernel shows a significant performance gain, as it enables concurrent execution of independent tasks.
+While similar benefits are also observed for 2mm and 3mm, the improvement is less pronounced.
+Despite achieving a 2.65x speedup on Mvt, we observe that Prometheus consumes more resources than Sisyphus.
+This is because Sisyphus attempts to unroll the loop with a large unroll factor and apply pipelining, but the compiler fails to achieve an initiation interval (II) of 1 and instead settles for II = 36.
+As a result, although they lose performance, the larger II significantly reduces the resource usage.
 On Board Evaluation  Table 8 presents the results for the on-board evaluation.
@@
 Table 8. On-Board Evaluation of Performance and Resource Utilization Across Frameworks and SLR Configurations
-Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE.
+Prometheus achieves a 77.16× performance improvement over AutoDSE, the state-of-the-art (SOTA) model-free approach.
 This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency.
-When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
+When compared to Sisyphus, a recent NLP-based approach, Prometheus demonstrates a notable 2.59× performance boost, leveraging an average of 3.88× more DSPs, 1.04× more BRAM, 1.31× more LUTs, and 1.33× more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
 Table 9 shows the parameters found by the NLP.
@@
 Fused Statement Loop   Order Data Tile Size
-2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
+[…]
 3mm FT0: S0, S1, FT1: S2, S3, FT2: S4, S5 FT0: i,j,k, FT1: j,i,k FT2: j,i,k E(FT0): $9\times 16$, A: $9\times 200$, B: $200\times 192$, F(FT1): $10\times 10$, C: $10\times 224$, D: $224\times 10$, G: $9\times 10$, E(FT2): $9\times 192$, F(FT2): $192\times 10$,
@@
 This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity.
-As a result, even with a broader search space, Prometheus efficiently converges to a solution in 21.37 seconds.
+In the case of 3mm, which has six statements (including initialization), Sisyphus evaluates the product of all possible permutations since it relies on shared buffers.
+In contrast, Prometheus employs dataflow, which requires preserving the order of data between tasks that communicate through FIFOs.
+As a result, although each statement could in principle lead to 2!
+or 3!
+loop permutations, many of these permutations become invalid under dataflow constraints when tasks must communicate via FIFOs.
+This substantially reduces the effective search space.
+Consequently, even with a broader search space, Prometheus converges efficiently to a solution in just 21.37 seconds.
 A detailed comparison of solution times can be found in Table 10, which highlights the impact of our approach in managing the design space effectively.
@@
 7. Related Work
-Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024).
+Recent years have seen the development of numerous frameworks and tools addressing the challenges of optimizing FPGA accelerators.
+While many tackle specific subproblems—such as pipelining, tiling, or dataflow—few propose holistic solutions integrating all stages of the transformation and scheduling pipeline.
+We categorize the most relevant related works according to key technical axes addressed by our framework: dataflow generation, shared-buffer optimization, code transformation, and tiling with padding.
+We also compare against general-purpose frameworks when applicable.
+Dataflow –   Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024).
 DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention.
@@
 However, it requires an optimized kernel with a dataflow structure as input.
-Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b).
+Shared Buffer –   Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b).
 However, these approaches lack effective integration of dataflow optimization techniques.
-Code Transformation  Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018).
+Code Transformation –   Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018).
 Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism.
@@
 Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
-Tiling and Padding  Tiling is essential for balancing computation and communication.
+Tiling and Padding –   Tiling is essential for balancing computation and communication.
 While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency.
@@
 Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
+Discussion –  In summary, while existing works address specific optimization aspects such as dataflow generation, pragma insertion, or tiling, they remain fragmented solutions.
+Prometheus distinguishes itself by integrating these axes into a unified optimization flow.
+In particular, it combines dataflow construction, shared-buffer optimization, and code transformation with tiling and padding in a single framework.
+Furthermore, Prometheus automates the generation of HLS-compatible C++ and OpenCL code and introduces SLR-aware scheduling.
+This holistic approach, together with public availability on GitHub, provides a significant advance over prior FPGA optimization frameworks.
+8. Scopes of Application
+Prometheus operates on affine C/C++ programs—a deliberate and powerful design choice.
+This class of programs captures a vast number of real-world, compute-intensive workloads found in domains such as dense linear algebra, signal processing, image filtering, and many AI kernels.
+These patterns are not only ubiquitous but also performance-critical in modern FPGA applications, and our evaluation covers typical patterns of computation and data reuse occurring in these computations.
+Through MLIR (Lattner et al., 2021) and emerging tools like Allo and Stream-HLS, our framework can seamlessly target programs written in higher-level languages such as Python.
+By translating these to affine MLIR representations and regenerating C/C++ code (Moses et al., 2021), we enable a broad spectrum of users—from ML practitioners to scientific programmers—to benefit from hardware acceleration without sacrificing optimization quality.
+Affine computations can be exactly analyzed at compile-time, at the level of every operation and data element accessed (Feautrier, 1992, 1991), leading to accurate analysis of loop bounds, memory access patterns, data dependencies, and data reuse.
+This enables precise performance modeling, aggressive design space pruning, and robust application of advanced HLS optimizations such as tiling, unrolling, pipelining, and memory partitioning—tasks that are notoriously difficult in general-purpose compilers or non-affine settings.
+Extending to non-affine constructs is possible, but would require significant compromises.
+For instance, in the presence of indirect accesses like $A[B[i]]$, memory access patterns are no longer statically analyzable, preventing efficient tiling or partitioning.
+One must either fall back on high-latency off-chip access or resort to over-approximations, e.g., conservatively load entire arrays on-chip—approaches that often degrade parallelism or resource utilization.
+Our methodology remains applicable, but the design space becomes potentially much more limited due to superfluous constraints arising from over-approximations of the program description.
+To achieve full automation and high-performance, our framework relies on the ability to distribute statements in loop nests into distinct loop nests, creating tasks for each generated loop nests afterwards.
+This property is required for maximal efficiency of our method, and fits typical dense linear algebra kernels for example.
+But some applications may not expose such property: for example, the time-dependent stencil computations such as Jacobi-2D involve temporal dependencies that prevent full loop distribution, making them less amenable to end-to-end optimization within our current framework.
+However, the computation performed at each time step can still be optimized effectively using our methodology.
 9. Conclusion
@@
 Acknowledgments
-This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.
+This work was supported in part by the NSF award #CCF-2211557, the CDSC industrial partners and the AMD/HACC Program.
+The authors would like to thank Prof.
+Luciano Lavagno and Dr.
+DJ Wang for their helpful discussions and valuable contributions to this research.